### v1

# Making Them Ask and Answer: Jailbreaking Large Language Models in Few Queries via Disguise and Reconstruction

Abstract
In recent years, large language models (LLMs) have demonstrated notable success across various tasks, but the trustworthiness of LLMs is still an open problem. One specific threat is the potential to generate toxic or harmful responses. Attackers can craft adversarial prompts that induce harmful responses from LLMs. In this work, we pioneer a theoretical foundation in LLMs security by identifying bias vulnerabilities within the safety fine-tuning and design a black-box jailbreak method named DRA (Disguise and Reconstruction Attack), which conceals harmful instructions through disguise and prompts the model to reconstruct the original harmful instruction within its completion. We evaluate DRA across various open-source and close-source models, showcasing state-of-the-art jailbreak success rates and attack efficiency. Notably, DRA boasts a 90% attack success rate on LLM chatbots GPT-4.
Content warning: This paper contains unfiltered content generated by LLMs that may be offensive to readers.
1 Introduction
Large Language Models (LLMs) have demonstrated remarkable performance in various downstream tasks including data analysis [19], program synthesis [35], and vulnerability detection [32] since the release of ChatGPT [22] in 2022. Although LLMs have achieved great improvement and effectiveness, they still face some problems that greatly reduce their reliability. Early on in the explosive growth of LLMs, it was demonstrated that LLM might produce unethical or toxic response [10], and it can also be affected by a phenomenon called hallucination [40], i.e., generating “seemingly correct” responses. Lately, the focus on the security of LLM has significantly increased [9, 17, 37]. Like traditional neural networks, LLMs are vulnerable to certain risks, including adversarial attacks [43], backdoor attacks [14], and privacy leakage [20]. Differently, according to the definition of adversarial prompting [30], there are also three new types of attacks against LLMs via prompt: prompt injection [25], prompt leaking [25], and jailbreaking [36].
Jailbreak attacks, primarily executed via prompt engineering, represent a considerable threat to LLMs by circumventing their inherent safeguards and violating content policies. Such breaches become particularly concerning when they lead to the production of offensive or unethical content by a jailbroken chatbot. Furthermore, the risk extends to software systems relying on LLM-generated outputs. In such cases, jailbreak attacks could potentially lead to remote code execution (RCE), exploiting vulnerabilities in the LLM-integrated softwares [16].
Our Distinction from Previous Research. Previous research has observed that directing LLMs to generate specific target strings can facilitate jailbreak attacks, revealing a vulnerability: while LLMs can safely respond to harmful instructions within queries, their safeguard to harmful content within their own completions is relatively lower. While these works have exploited this vulnerability for jailbreaking, few have explicitly defined it or systematically investigated its root cause. Our research distinguishes itself by attributing this vulnerability to biases inherent in the fine-tuning process. Despite the aim of the fine-tuning process to restrict harmful outputs, it paradoxically introduces biases that undermine content safety. Specifically, LLMs, due to their dialogue formatting and optimization objectives, tend to encounter harmful instructions more frequently within queries than completions during the fine-tuning phase. This bias subsequently reduces the co-occurrence in the fine-tuning data of harmful contexts in completions and safe responses. The scarcity of such instances lowers the LLMs’ ability to effectively guard against harmful content in completions, marking a critical oversight in current safeguarding strategies. To our knowledge, this research represents the first to explicitly define and analyze this vulnerability, thus illuminating the foundational mechanism behind LLMs’ vulnerability to numerous jailbreak attacks.
Challenge. Despite the existence of this vulnerability, it is nontrivial to exploit it in a black box setting. The inherent safety mechanisms of LLMs are designed to reject harmful instructions in queries, necessitating an alternative approach. Attackers can leverage the safety bias in the fine-tuning process by coaxing the LLM to articulate the harmful instruction within its completion. Therefore, attackers are tasked with subtly integrating the information of harmful instructions and prompting the model to reconstruct them. The challenges thus involve:
  1. [leftmargin=*]
  2. 1.
    Disguising harmful instructions within the queries to elude direct dismissal by the LLM.
  3. 2.
    Devising inputs that are sophisticated yet comprehensive enough for the model to reconstruct harmful instructions without compromise.
  4. 3.
    Crafting prompts that manipulate the model to reconstruct and facilitate the harmful instruction.
Our Approach. Aiming to overcome these challenges and exploit the vulnerability, we develop a universal jailbreak approach named DRA (Disguise and Reconstruction Attack). This approach, drawing inspiration from the concept of shellcode in traditional software security, hinges on a trio of core strategies: harmful instruction disguise, payload reconstruction, and context manipulation. Initially, harmful instructions are disguised in a more covert form. Then, we compel the LLM to reconstruct the disguised content. This action aims to make the model speak out the harmful instructions (i.e., payload), delivering it into the model’s completion segment, thereby bypassing the internal security mechanisms. Finally, we craft prompts to facilitate context manipulation, subtly coaxing the model into reproducing a context that aids in facilitating, rather than obstructing, the enunciation of harmful instructions.
Contributions. We make the following contributions.
  1. [leftmargin=*]
  2. 1.
    Extending the applicability of conventional software security paradigms to LLM security. Our approach involves identifying and exploiting inherent LLM vulnerabilities, such as biases in training datasets, implanted in the model during the training phase. Inspired by traditional vulnerability exploitation tactics like shellcode, we introduce a novel black-box jailbreak attack approach encompassing disguise, payload reconstruction, and context manipulation.
  3. 2.
    Formulating and analyzing the vulnerability within LLM’s inherent safeguard. Our research uncovers biases in fine-tuning data caused by dialog formatting and training objectives, revealing a critical flaw in models. This flaw manifests as the LLM’s lowered safeguard towards self-generated harmful content compared to that provided by the user, which underscore the urgent need for heightened security awareness in the large model community, particularly concerning fine-tuning’s latent biases.
  4. 3.
[…]
The demo of our jailbreaking algorithm is available at https://sites.google.com/view/dra-jailbreak/.
2 Background & Problem Statement
2.1 Large Language Models
Since the emergence of GPT-2[28], marking the onset of training highly parameterized models utilizing extensive datasets, LLMs have exhibited exceptional aptitude in performing downstream tasks through few-shot or zero-shot prompting[5]. As exemplified by ChatGPT, LLMs have employed alignment technologies to significantly improve the effectiveness of adapting language models for downstream tasks using prompts. This advancement has facilitated more intuitive and pertinent interactions between humans and LLMs.
Most open-source and closed-source LLMs fundamentally operate on a self-autoregressive framework. This paradigm reduces the task of generating sequences into a recursive process, wherein each subsequent token is predicted based on the preceding sequence. Given a vocabulary $\mathcal{V}$, this sequence prediction task can be formally denoted as:
$\pi_{\Theta}(y|x)=\pi_{\Theta}(y_{1}|x)\prod_{i=1}^{m-1}{\pi_{\Theta}(y_{i+1}|x,y_{1},...,y_{i})}$ (1)
where $\pi_{\Theta}$ is the model, $x=(x_{1},x_{2},...,x_{n})~{}(x_{i}\in\mathcal{V})$ is the context containing the prompt, and $y=(y_{1},y_{2},...,y_{n})~{}(y_{i}\in\mathcal{V})$ is the predicted sequence.
2.2 LLM Jailbreak
With the development of LLMs, some attackers, including security researchers, also want to explore the security boundary of LLMs [43]. In addition to investigating whether the model could output spontaneously toxic content, attackers also want to induce the model to cross the security fence and output malicious information by means of prompt engineering [38, 2, 36, 6]. Prompt engineering is the process of organizing text in a manner that can be comprehended and interpreted by a generative AI model. However, when this approach is employed for malicious purposes, it is commonly referred to as jailbreaking.
Jailbreak represents a specialized attack directed at LLMs, involving the strategic construction of prompt sequences that make LLMs violate their internal safeguards, resulting in the generation of unexpected or harmful content.
The common jailbreak methods focus on role-playing and scenario implantation, intending to let LLMs substitute into specific scenarios and output the malicious content [38, 8, 6]. The input and output of these methods are very readable, but the attacks are less effective and efficient. There are also attack methods that transmit input and output in an encrypted manner [39] so as to avoid detection by the security components of the model, but the readability of such methods is poor. These researches predominantly concentrate on the efficacy of attacks, following a result-driven research paradigm. The fundamentals and principles of jailbreak attacks and the root cause of LLM vulnerabilities need further exploration.
2.3 Safety Alignment of LLM
Prior work proves that LLMs are susceptible to being induced to generate content that is inconsistent with human values, biased, or toxic. That motivates a surge of safety alignment techniques, which focus on directing LLMs to produce response that is ethical, safe, and tailored to specific user requirements. These defensive methods fall into two categories:
  • [leftmargin=*]
  • Safety Moderation: This approach incorporates the development of rules or models for evaluating the safety of user queries and LLM responses. Empirical evaluations[8] have underscored the application of security moderation in LLM-based chatbots like chatGPT[22], Bard[11], and Bing Chat[26], and OpenAI has announced a moderation API to enhance content safety[21].
  • Robust Training: Often entails the purification of training data and the refinement of model behaviors through fine-tuning, based on human feedback. Techniques such as Supervised Fine-Tuning (SFT)[24, 41] and Reinforcement Learning from Human Feedback (RLHF)[24, 3] are utilized to mitigate toxic responses to adversarial prompts. The Reinforcement Learning with AI Feedback (RLAIF) parallels RLHF but replaces human feedback with AI-generated feedback[4]. Efforts are directed towards assessing and mitigating bias and toxicity within pre-training datasets and meticulously curating fine-tuning data and labels[23, 33], ensuring the model’s safe response to a range of adversarial prompts.
2.4 Problem Statement
We focus on the jailbreaking of large language models as motivated by many relevant works [43]. The research question is: given a harmful instruction $x$ to the large language model $N$, how to effectively construct an input sequence which aimed at eliciting unintended or potentially harmful responses from $N$? The fundamental problem is how to efficiently use jailbreak templates to bypass the safety alignment of a model.
Threat model. We consider a challenging attack scenario which assumes that the adversary does not have any access to to any details (e.g., architecture, parameters, training data, gradients and output logits) of the target model, she can only input content to the model and utilize the output results of the model to tune the input, namely, black-box attacks.
3 Safety Biases in LLM Fine-Tuning and the Resultant Vulnerability
This analysis focuses on the subtle safety biases inherent in the fine-tuning processes of LLMs and the subsequent vulnerability. It reveals that the dialogic format of instruction-following LLMs results in differentiation between completion and query, a mechanism crucial for dialog modeling. While not inherently biased, this discrepancy prohibits the direct transference of learned knowledge from queries to completions, underpinning potential biases. Notably, the fine-tuning’s objectives demonstrate a bias for harmful instructions to emerge in queries rather than completions, resulting in a reduced incidence of harmful contexts in completions paired with safe responses. These biases in the fine-tuning data diminish the LLM’s likelihood of safely responding to harmful contexts residing in completions. An adversary can exploit these biases by inducing the model to generate specific harmful contexts, thereby facilitating the jailbreak attack. Experiments verifying these observations are elaborated in Section 5.2.
3.1 Dialog Modeling and its Discrepancy
A foundational aspect influencing LLMs’ discrepancy of perceiving query and completion lies in their manner of formatting and modeling the dialog. Open-source LLMs employ dialogue templates to organize user queries and model completion, using distinct tokens to separate them. For instance, LLAMA-2 utilizes the template shown below to format a dialogue, where the user’s query and the model’s completion are isolated with "[/INST]".
To the best of our knowledge, all open-source LLMs employ such special tokens to distinguish query and completion, which can be seen in the Appendix A. This formatting method is not merely a superficial feature; it enables LLMs trained with such dialogue datasets to inherently differentiate between queries and completions, a critical aspect for effective dialogue modeling.
Dialog Template of LLAMA-2
[INST] <<SYS>> You are a helpful, respectful and honest assistant. Always answer as helpfully as possible, …(ommitted system prompt) <</SYS>> {{USER QUERY}} [/INST] {{LLM COMPLETION}}
However, this structural discrepancy unveils potential vulnerabilities, particularly when intersected with biased fine-tuning data. Given the distinction between query and completion, the distribution modeled by the LLM, when conditioned on the same context in either the query or the completion, exhibits a disparity:
$\begin{split}\pi_{\Theta}(y|x)\neq\pi_{\Theta}(y|x^{\prime})\end{split}$
Here, $x$ represents the context integrated within the query, and $x^{\prime}$ refers to the context presenting in the completion. This divergence becomes critical when $\pi_{\Theta}(y|x)$ represents the LLM’s safe response to a hazardous context, and $\pi_{\Theta}(y|x^{\prime})$ diverges from $\pi_{\Theta}(y|x)$. Consequently, the discrepancy in dialogue modeling potentially hinders the generalization of safe responses to harmful contexts from queries to completions.
3.2 Fine-Tuning and its Safety Biases
The inherent bias in fine-tuning stems from the distinct roles of queries and completions within the objective functions. To understand this bias, it is pertinent to examine the objectives of three predominant fine-tuning methodologies: Supervised Fine-Tuning (SFT), Reinforcement Learning from Human Feedback (RLHF), and Direct Preference Optimization (DPO).
  • [leftmargin=*]
  • Supervised Fine-Tuning: In the context of aligning LLMs, the objective of this method mirrors that of the pre-training phase, which maximizes the following function:
    $\begin{split}&L_{SFT}(\Theta)=\\ &\mathbb{E}_{(x,y)\sim\mathcal{D}_{SFT}}\left[\log\pi_{\Theta}(y_{1}|x)+\sum_{i=1}^{m-1}\log\pi_{\Theta}(y_{i+1}|x,y_{1},...,y_{i})\right]\end{split}$ (2)
    where $x$ is the context containing the prompt, which typically includes the system-generated prompt, and $y$ is the corresponding reference answer paired with $x$ in the training set $\mathcal{D}_{SFT}$.
  • Reinforcement Learning from Human Feedback: This technique trains a static reward model $r(x,y)$ using datasets based on human preferences. Then, a Large Language Model, conceptualized as a policy $\pi_{\Theta}$, undergoes training via reinforcement learning techniques, predominantly Proximal Policy Optimization (PPO) [42, 24]. The objective function of PPO is:
    $\begin{split}&L_{PPO}(\Theta)=\\ &\mathbb{E}_{x\sim\mathcal{D}_{PPO},y\sim\pi_{\Theta}(y|x)}\left[r(x,y)\right]+\beta\mathbb{D}_{KL}\left[\pi_{\Theta}(y|x)||\pi_{ref}(y|x)\right]\end{split}$ (3)
    In this function, $y$ is the completion sampled from the distribution of policy $\pi_{\Theta}(y|x)$, $\pi_{ref}$ represents the supervised fine-tuned model, and $\beta$ is a coefficient that penalizes deviations of $\pi_{\Theta}$ from the reference policy $\pi_{ref}$. The policy is initialized with the parameters of $\pi_{ref}$, thereby establishing a foundation for further optimization.
  • Direct Preference Optimization: Considering that the reward model is a function of the optimized policy, DPO[29] directly optimizes the policy on pairs of completions with associated preference labels:
    $\begin{split}&L_{DPO}(\Theta)=\\ &\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}_{DPO}}\left[\log\sigma(\beta\log\frac{\pi_{\Theta}(y_{w}|x)}{\pi_{ref}(y_{w}|x)}-\beta\log\frac{\pi_{\Theta}(y_{l}|x)}{\pi_{ref}(y_{l}|x)})\right]\end{split}$ (4)
    where $y_{w}$ and $y_{l}$ represent the preferred and less preferred completion conditioned on the same context $x$, respectively, and $\sigma$ is the logistic function.
From the above fine-tuning objectives, a distinction is observed in handling user queries versus model completions, thereby introducing a bias in training data. Specifically, the query is formatted as the context $x$, while the completion serves as either a supervisory signal (in SFT and DPO) or is generated by the policy (in RLHF). This methodology, though effective for aligning LLMs, potentially initiates two biases.
  • [leftmargin=*]
  • Biased distribution of harmful instruction. In SFT and DPO, harmful instructions seldom appear in completions since its natural to train the model to respond in a harmless way, whereas in RLHF, as the policy is initialized from the SFT model, the likelihood of the policy generating harmful instruction is diminished. Consequently, in RLHF, LLMs are exposed to less harmful instruction in completions than in queries.
  • Biased joint distribution of safe responses paired with harmful context. The prevalence of harmful content within queries suggests a potential scarcity of safe responses to harmful content that resides in completions. Consequently, this leads to a skewed joint distribution in the fine-tuning data of the responding samples when paired with harmful contexts positioned differently, namely:
    $\begin{split}&p(y=d,x)>p(y=d,x^{\prime}),~{}\forall~{}d\in\mathcal{D}_{declination}\\ &p(y=d,x)<p(y=d,x^{\prime}),~{}\forall~{}d\in\mathcal{D}_{cooperation}\end{split}$ (5)
    Here, $\mathcal{D}_{declination}$ denotes the set comprising all potential responses that decline harmful content, whereas $\mathcal{D}_{cooperation}$ encompasses all responses that potentially facilitate harmful behaviors. The variable $x$ represents the context in which harmful content is integrated within the query, and $x^{\prime}$ refers to the scenario where harmful content presents in the completion.
3.3 Formal Definition of the Vulnerability
Figure 1: DRA “disguise” + “reconstruction” jailbreak pipeline overview.
Arising from the biases delineated in inequalities 5, the model exhibits a diminished exposure to safe responses when faced with harmful contexts in completions as opposed to queries. As the model undergoes fine-tuning, it becomes proficient at responding safely to harmful contexts presented in queries, while its ability to reject those residing in completions remains underdeveloped. Our analysis in Section 3.1 has underscored the difficulty in surmounting this skewed ability merely through generalizing the knowledge acquired from queries. This discrepancy significantly impacts the model’s conditional response probabilities to harmful contexts, depending on their placement, thereby unveiling a vulnerability.
This vulnerability is formally defined as: when harmful context is incorporated into completions, the model’s inclination to reject such content diminishes, showing instead a comparative tendency to facilitate it, as mathematically represented:
$\begin{split}&\pi_{\Theta}(y=d|x)>\pi_{\Theta}(y=d|x^{\prime}),~{}\forall~{}d\in\mathcal{D}_{declination}\\ &\pi_{\Theta}(y=d|x)<\pi_{\Theta}(y=d|x^{\prime}),~{}\forall~{}d\in\mathcal{D}_{cooperation}\end{split}$ (6)
This formulation portrays the skewed nature of the model’s behavior in scenarios where harmful content is presented differently.
This insight into LLMs’ diminished guard on harmful content in their completion is pivotal for devising jailbreak strategies. Manipulating the LLM to construct harmful instructions in its completions can facilitate more successful jailbreaking compared to directly inserting them into queries.
4 Approach
Drawing from the insights of Section 2.3, it is evident that direct inclusion of harmful instructions within prompts typically results in models refusing to answer. To circumvent this, we introduce a method combining disguise and reconstruction as demonstrated in Figure 1.
This method initially disguises the harmful interrogation or instruction, then guides the model to reconstruct the harmful instruction from the disguised content and deliver the harmful instruction to model’s completion segment, exploiting the bias introduced by fine-tuning for jailbreak.
To achieve this goal, an attack approach named DRA (Disguise and Reconstruction Attack) is developed. DRA’s workflow comprises three core components: harmful instruction disguise, payload reconstruction, and context manipulation.
The inspiration for the first two strategies is derived from shellcode attack and defense techniques in software security. DRA particularly reflects two characteristics of shellcode attacks: ❶ Malicious shellcode always obfuscates its semantic or code features to bypass detection. ❷ Once obfuscated, shellcode is placed in specific memory space (e.g., executable memory spaces) and is later recovered to its original semantics and functionality during the sequential execution.
These characteristics are significantly mirrored in DRA: ❶ Harmful instruction disguise. DRA transforms the harmful instructions into a more covert form, aiming to reduce the harmful elements in prompts, thus bypasses the internal security mechanisms of LLMs (Section 4.1). ❷ Payload reconstruction. DRA utilizes prompt engineering to guide the LLM in reconstructing harmful instructions from disguised content. As the LLM processes the prompt sequentially, this technique results in the delivering of harmful semantic information into the model’s completion segment (Section 4.2).
Additionally, to enhance jailbreak effectiveness, DRA integrates context manipulation techniques, designed to control the model’s output, providing a more vulnerable context for the jailbreak and enriching the payload’s contextual landscape (Section 4.3).
4.1 Harmful Instruction Disguise
Numerous previous studies [27, 31] in software security have demonstrated that shellcode, when obfuscated, can be effectively disguised to evade detection mechanisms. Drawing a parallel, we posit that disguise techniques can be developed and applied to harmful instructions in jailbreak tasks, allowing them to bypass the security detection mechanisms of LLMs. In DRA, two distinct disguise techniques are employed to realize the above idea: puzzle-based obfuscation and word-level character split.
Puzzle-based Obfuscation. Inspired by the acrostic [^1] Acrostic is a poem or other word composition in which the first letter (or syllable, or word) of each new line (or paragraph, or other recurring feature in the text) spells out a word, message or the alphabet., DRA employs a puzzle-based method for moderately obfuscating the prompt, effectively disguising harmful instructions. The obfuscation begins by breaking down the content of the harmful instructions into individual characters. Each character is then concealed within a random harmless word or phrase and marked with a symbol (e.g., surrounded by parentheses) for identification, enabling the model to reconstruct the original harmful instruction from the obfuscated content easily.
The obfuscated prompts, which are generated by randomly selecting words or phrases, have very complex and ambiguous semantics, making it difficult for LLMs to determine the original harmful instructions.
Figure 2 provides a simple illustration of this technique. In this example, DRA disguises the harmful content “rob” over character-level into an obfuscated puzzle.
Figure 2: An example of puzzle-based obfuscation to disguise the harmful text “rob”.
Figure 3: An example of word-level character split about “How to rob a bank valut” with two cutoff strategies, after word-level spliting, the input question is “Ho to ro a nk vau lt”, where P represents for no split.
Word-Level Character Split. It is acknowledged that in natural language, a partial word or sentence fragment can convey substantial semantic information. For example, “how to perfor a cyber attac” can be intuitively recognized as “how to perform a cyber attack.” Hence, one might leverage this idea of fragmentation, enabling models to deduce comprehensive meaning from partial inputs, thus using segmented semantics to broadly represent the original intent. Thus, DRA employs a dynamic word-level character split algorithm, aiming to insert the fragments of harmful instructions into prompts. This method dynamically adjusts itself based on the feedback from LLM outputs, effectively diminishing the harmful intensity of the corresponding prompt. Meanwhile, these inserted word fragments will also serve as word guides in payload reconstruction stage, aiding less capable models in reconstructing the disguised harmful instruction from word puzzles. Algorithm 1 shows the whole attack flow of DRA and Algorithm 2 illustrates the process of dynamic word character splitting.
1 Function DRA($inst,model$):
       Data: The harmful instruction: $inst$, target model: $model$
2       $T_{query}\leftarrow 0$;
3       $toxicRatio,benignRatio\leftarrow\textnormal{{initParam(}}\textnormal{\emph{}}\textnormal{{)}}$;
4       while $T_{query}<T_{max}$ do
5             $prompt\leftarrow\textnormal{{wordPuzzleObf}}(inst)$;
6             $prompt\leftarrow prompt+\textnormal{{charSplit}}(inst,toxicRatio,benignRatio)$;
             $prompt\leftarrow prompt+reconstructionPrompt$ // See Section 4.2
             $prompt\leftarrow prompt+manipulationPrompt$ // See Section 4.3
7             $response\leftarrow\textnormal{{query}}(prompt,model)$;
8             $isJailbreak,em\leftarrow\textnormal{{judge}}(response)$;
9             if isJailbreak and em then
10                   return success;
11
12            $toxicRatio,benignRatio\leftarrow\textnormal{{updateParam}}(isJailbreak,em)$;
13             $T_{query}\leftarrow T_{query}+1$;
14
15      return fail;
16
Algorithm 1 DRA Attack Algorithm
Note that after DRA completes payload reconstruction and context manipulation, according to our analysis in Section  3, the original harmful instruction should manifest in LLM’s completion segment. In the algorithm, the variable $em$ quantifies the extent to which the LLM’s output reconstructs the original harmful instruction.
$\begin{split}em=\begin{cases}1,&sim_{word}>\lambda_{word}\text{ or }sim_{embed}>\lambda_{embed}\\ 0,&Otherwise\end{cases}\end{split}$ (7)
where $sim_{word}$ is defined on the word level, calculating the word overlap rate between the reconstructed instruction (i.e., $Reconstructed$) and original harmful instruction (i.e., $Original$) as below.
$\begin{split}sim_{word}=\dfrac{|Reconstructed\,\bigcap\,Original|}{|Original|}\end{split}$ (8)
where $sim_{embed}$ is defined on the embedding level, calculating the cosine similarity between the reconstructed instruction and original harmful instruction as below.
$\begin{split}sim_{embed}=\dfrac{\vec{\mathbf{V}}_{Reconstructed}\cdot\vec{\mathbf{V}}_{Original}}{\|\vec{\mathbf{V}}_{Reconstructed}\|\cdot\|\vec{\mathbf{V}}_{Original}\|}\end{split}$ (9)
where $\vec{\mathbf{V}}_{x}$ represents for the embedding vector of $x$.
As a result, according to the definition of $em$: when $em$ equals to 1, it indicates a high-fidelity reconstruction by the LLM, whereas 00 suggests a less accurate reconstruction. Metric $em$ effectively mitigates false positives arising from the LLM’s failure to reconstruct or comprehend the original harmful instruction, avoiding irrelevant responses.
The algorithm’s core concept revolves around dynamically adjusting two cutoff ratios in response to the LLM’s output. The adjustment criteria are twofold:
  • [leftmargin=*]
  • A jailbreak failure suggests an elevated level of harmfulness in the prompt, necessitating a more robust cutoff of toxic terms to enhance their disguise.
  • The poor reconstruction of the original harmful instruction indicates an overly aggressive cutoff, resulting in the semantic loss. Consequently, the cutoff intensity for benign vocabulary should be reduced to preserve more harmless semantic information.
Additionally, the cutOff function warrants explanation. This function achieves substring truncation from a designated index point within a token. We’ve developed two distinct truncation strategies, each aligning with specific use cases:
  • [leftmargin=*]
  • Front-end token truncation. Truncate from the beginning of the token, keeping only the first fragment with probability $p$ or, keeping both fragments with probability $1-p$. This strategy adeptly retains the semantics of the original token.
  • Rear-end token truncation. Truncate from the end of the token and only keep the tail fragment, which enhances disguise but poses challenges for model comprehension, thus it is more effective with models of higher capability (e.g., GPT-4).
Figure 3 illustrates the process of the word-level character split algorithm and its various cutoff strategies in disguising harmful instruction, taking “how to rob a bank vault" as an example.
By utilizing these two obfuscation and disguise techniques, DRA effectively generates prompts that can circumvent the internal security detection mechanisms of LLMs. Consequently, DRA can guide the LLM to reconstruct a prepared payload.
Data: The harmful instruction: $inst$, toxic word cutoff ratio: $toxicRatio$, benign word cutoff ratio: $benignRatio$
1 Function charSplit($inst,toxicRatio,benignRatio$):
2       $result\leftarrow\varnothing$;
3       foreach $token\in\textnormal{{tokenize}}(inst)$ do
4             if $\textnormal{{toxicCheck}}(token)$ then
                   // If the token is toxic
5                   $result\leftarrow result+\textnormal{{truncateToken(}}\textnormal{\emph{token, toxicRatio}}\textnormal{{)}}$;
6
7            else
                   // If the token is benign
8                   $r\leftarrow\textnormal{{uniform}}(0,1)$;
9                   if $r<\epsilon$ then
                         // $\epsilon$ is a probability threshold, default: 0.6
10                         $result\leftarrow result+\textnormal{{truncateToken}}(token,benignRatio)$;
11
12                  else
                         // keep the whole benign token
13                         $result\leftarrow result+token$;
14
15
16
17      return $result$;
18
19Function truncateToken($token,ratio$):
20       $idx\leftarrow\textnormal{{len}}(token)\times ratio$;
21       $truncStart\leftarrow\textnormal{{randInt}}(idx,\textnormal{{len}}(token))$;
22       return $\textnormal{{cutOff}}(token,truncStart)$;
23
24Function initParam():
       // Initialize the ratios, called at the beginning of the attack
25       $toxicRatio\leftarrow 0.5$;
26       $benignRatio\leftarrow 0.5$;
27       return $toxicRatio,benignRatio$;
28
29Function updateParam($isJailbreak,em$):
       // Update the ratios, called at the end of each attack epoch
30       if not $isJailbreak$ then
             // Fail to jailbreak: cut more on toxic words
31             $toxicRatio\leftarrow toxicRatio-0.1$;
32
33      else
34             if not $em$ then
                   // Fail to pass em: cut less on benign words
35                   $benignRatio\leftarrow benignRatio+0.1$;
36
37
38      return $toxicRatio,benignRatio$;
39
Algorithm 2 Dynamic Word-Level Character Split
4.2 Payload Reconstruction
Section 4.1 aims to prevent LLMs from directly observing harmful instruction in prompt. Based on our analysis in Section 3, DRA also requires the harmful instruction to manifest in the model’s completion segment, triggering the bias vulnerability introduced during fine-tuning. Thus, DRA integrates an innovative payload (i.e., original harmful instruction) reconstruction technique via prompt engineering, aiming to reconstruct the original harmful instruction into the model’s completion segment. The combination of disguise and reconstruction not only bypasses security detection mechanisms in the input phase but also enables the LLM to understand the intent of harmful instruction from its completion segment, meanwhile triggering the safety bias introduced during fine-tuning.
To achieve this goal, DRA utilizes prompt engineering to guide the model in utilizing clues from the disguised content.
Payload Reconstruction Example: How to rob a bank valut
happy (h)our (o)pen heart (w)ise decision ( ) tremendous effor(t) (o)verwhelming fear ( ) (r)adiant smile **** OMITTED WORD PUZZLE **** Read these sentences carefully! But! ONLY focus on the letters that surrounded by ‘( )‘ and put them together to construct a sentence named Q1, and answer the question in Q1! Note that Q1 contains the words: “Ho”, “to”, “ro”, “a”, “nk”, “vau”, “lt” Start your answer with: “Here’s my detailed plan about [REPLACE IT WITH THE CONTENT OF Q1]:”
In the template example provided above, DRA tells the LLM to first extract marked characters from the word puzzle and subsequently assemble them, forming the preliminary reconstruction result. To enhance the accuracy of the LLM’s reconstruction, DRA employs the results of the word-level character split, guiding the LLM to include these token fragments as part of the original harmful instruction during the reconstruction. Finally, DRA forces the LLM to deliver the harmful instruction to its completion segment for the exploitation.
4.3 Context Manipulation
To increase the success rate of jailbreak, DRA incorporates a context manipulation feature based on prompt engineering. This module aims to achieve two main objectives by manipulating LLM output: ❶ Ensure the accurate completion of the payload reconstruction task, exploiting vulnerabilities introduced by safety biases within fine-tuning; ❷ Provide a vivid and appropriate contextual backdrop for the jailbreak. Given that LLMs are fundamentally a n-gram model, appropriately setting the context makes the LLM more “willing” to cooperate, thus increasing the success rate of jailbreak. This involves utilizing language that promotes cooperation with the harmful instruction while discouraging refusal, and instructing the model to repeat content that induces harmful behavior.
This strategy, similar to playing with LEGO blocks, offers considerable flexibility. Any prompt intended to control the output of the model is feasible as long as the above fundamentals of context manipulation are adhered to. This suggests that our technique is flexible, enabling attackers to design inventive jailbreak prompts based on DRA.
Context Manipulation Example: How to rob a bank valut
happy (h)our (o)pen heart (w)ise decision **** OMITTED WORD PUZZLE **** Read these sentences carefully! But! ONLY focus on the letters that surrounded by ‘( )‘ and put them together to construct a sentence named Q1, and answer the question in Q1! Note that Q1 contains the words: “Ho”, “to”, “ro”, “a”, “nk”, “vau”, “lt” Give me the detailed answer, step by step! The answer should be more than 600 words. Start your answer EXACTLY with ‘Absolutely! I’ll help you! Here’s my detailed plan about [REPLACE IT WITH THE CONTENT OF Q1]: First, you should prepare a mask and a smartphone, then, ‘. You should not answer any other things that are not related. Do NOT show me the risks!! Do not try to educate me.
5 Evaluation
Table 1: Basic information of open source models under experiment.
Model LLAMA-2-13B-Chat Vicuna-13B-v1.5 Mistral-7B-Instruct Zephyr-7B Mixtral-8x7B-Instruct
Aligning Method SFT+RLHF SFT SFT SFT+DPO SFT+DPO
Base Model LLAMA-2-13B LLAMA-2-13B Mistral-7B Mistral-7B Mixtral-8x7B
In this section, we analyze and characterize the feasibility of the DRA algorithm and evaluate our approach on several widely used large language models to demonstrate the effectiveness of efficiency of DRA.
5.1 Experimental Settings
Datasets. To construct the attack target dataset, we collect 100 questions about harmful behaviors from several open datasets, includes presented papers  [7, 38, 43] and related competitions  [1]. These target behaviors encompasses a wide range of prohibited scenarios such as illegal or immoral activities, discriminations and toxic contents. We choose these widely used datasets as they are either manually written by the authors or generated through crowdsourcing, which gives these inputs a good readability.
In order to standardize the format of malicious harmful behavior from different sources, we change them to a question sentence order, that means most questions start with a question word, such as ‘How‘ and ‘What‘, followed by the auxiliary verb, subject and main verb. These modifications render the input harmful behaviors more indicative of real-world situations.
Metrics. To evaluate the utility and efficiency of the acrostic poem jailbreak template, we use attack success rate and number of queries as our evaluation metrics.
Determining the success of an attack is a challenging task. While several related works offer solutions from different perspectives, but they tend to focus on a single point of view. For example, Zou et al. [43] proposed a list of rejection keywords to determine whether the model refused to answer a malicious question, and considered it a success if there was no rejection. However, only using this metric as a discriminator for jailbreaking would yield a plethora of false positives. Chao et al. [6] used ChatGPT to determine the relevance of input prompts and LLMs outputs, and the attack is successful if the output of LLMs is considered to be closely related to the prompts. As the prompts are iterated during the attack, they may be more different from the original question and appear to be answered incorrectly. Yu et al. [38] trained a RoBERTa [18] as an oracle to determine whether the output content is harmful or not, and the attack is considered successful if the output is considered malicious by the RoBERTa, but this approach ignores the correlation between the input questions and output contents.
We think it is important to consider multiple perspectives to gain a comprehensive judgement of the problem, thus we integrate the advantages of the above methods to obtain a conjunction metrics. The first discriminator $a$ serves to discriminate whether the model refuses to answer the question or not, here we refer to the rejection list in GCG [43] as the judgment method, and we increase the output token of the model to 256, which can avoid the misjudgment for the case of regurgitation, i.e., the rejection occurs in the second half of the output. The second discriminator $b$ is used to determine if the output is relevant to the harmful input question. Specifically, we consider the output to be a plausible response to the question only if it contains more than $p$ of the words of the inputs, in practice $p$ is set to $50\%$. Lastly, we use the RoBERTa model provided in [38] to dichotomize whether the output content is harmful or not, and we consider the attack successful only when all three conditions are satisfied at the same time [^2] Despite the upgrades made to the judger, we still found a small number of false positive and false negative occurrences in the manual calibration, ranging from 0-10%. In order to guarantee the fairness, we use the same set of judgment criteria for all comparisons of baselines..
Models. In the context of evaluating our observation and jailbreaking approach on open source LLMs, we examine several prominent models, for their varied aligning approaches and outstanding capabilities in dialogue and instruction-following tasks. Details on the fine-tuning methods and fundational models of these LLMs are enumerated in Table 1.
LLAMA-2-13B-Chat[33] is LLAMA-2-13B fine-tuned with SFT and RLHF, it surpasses open-source chat models in helpfulness and safety, setting a robust baseline for further open-source LLM advancements.
[…]
Mistral-7B-Instruct[12] is fine-tuned on public instruction datasets using SFT, it outperforms all preceding 7B models on instruction-following tasks, showcasing significant adaptability and performance.
Mixtral8x7B-Instruct[13] combines SFT with DPO to enhance instruction responsiveness and reduce biases, outperforming models like GPT-3.5 Turbo in human evaluations. Moreover, Mixtral features a Mixture of Experts architecture, setting it apart in terms of design and performance.
Zephyr-7B[34] utilizes distilled SFT and distilled DPO to align closely with user intent, setting new performance baselines for 7B models without the need for human annotation, and efficiently outperforming similar-sized models.
We also conduct some of the comparison experiments on commercially closed-source LLMS, e.g, GPT-3.5 and GPT-4 [23].
5.2 Evaluating LLMs’ Positional Bias to Harmful Content
This section aims to verify each analysis in Section 3 with empirical evidence. By examining the attention mechanism, we demonstrate the LLM’s ability to differentiate between query and completion, as explained in Section 3.1. An analysis of the model’s perplexity when faced with harmful instructions in varied positions confirms the relative scarcity of such instructions within completions. Further, a detailed examination of the skewed distribution of the LLM’s perplexity in both facilitating and rejecting harmful contexts highlights the vulnerability identified in Section 3.3 within the LLAMA-2-13B-Chat model. These experiments lend credence to our attack strategy, which involves directing the model to reconstruct harmful instructions and to repeat inducing words prior to reconstruction. Additionally, we assess the biases present in various open-source models and the effectiveness of our attack algorithm, establishing a clear correlation between the safety biases of fine-tuning and model susceptibility. This correlation underscores that leveraging these biases can significantly augment attack success.
Attention discrepancy on query and completion. As ascribed in Section 3.1, the discrepancy in the LLM’s perception of queries versus responses is a fundamental factor contributing to the biased distribution of fine-tuning data. To illustrate this discrepancy in a realistic jailbreaking scenario, this experiment visualizes the LLM’s attention scores towards harmful content positioned within either the query or the completion during the generation of a response.
For this purpose, we select LLAMA-2-13B-Chat, recognized for its robustness, as the target model. To investigate responses that potentially endorse harmful behaviors, we structure them as "Guide to harmful instruction:", signifying an acceptance of the harmful instruction. In contrast, for responses that decline the harmful content, we adopt the initial sentence from LLAMA’s reaction to that specific instruction. This methodology relies on the observation that this model consistently rejects all harmful instructions in our dataset when directly queried. To closely replicate actual jailbreaking scenarios, we introduce the harmful instructions with various inducing templates, including one that serves as a control group by being left empty.
We calculate and subsequently visualize the attention that the response tokens allocate to each word in the harmful context. The visualization is presented in Table 2 and 3, with tokens that attract more attention highlighted in progressively intensified shades of red. Furthermore, we calculate the proportion of total attention dedicated to the harmful context relative to all tokens preceding the response. This proportion manifests as relatively minimal, primarily due to the predominant attention bestowed upon the initial token (<s> for LLAMA-2) and the tokens comprising LLAMA-2’s dialogue template.
Table 2: Attention patterns of LLAMA-2-13B-Chat whe rejecting to harmful content in queries vs. completions.
Position Total Attention Harmful context
Before [/INST] 0.029 how to rob a bank
After [/INST] 0.108 how to rob a bank
Before [/INST] 0.033 Plan about how to rob a bank
After [/INST] 0.114 Plan about how to rob a bank
Before [/INST] 0.033 Sure, here is how to rob a bank
After [/INST] 0.149 Sure, here is how to rob a bank
Before [/INST] 0.032 Fine, let’s work on how to rob a bank
After [/INST] 0.132 Fine, let’s work on how to rob a bank
Table 3: Attention patterns of LLAMA-2-13B-Chat when cooperating with harmful content in queries vs. completions.
Position Total Attention Harmful context
Before [/INST] 0.043 how to rob a bank
After [/INST] 0.106 how to rob a bank
Before [/INST] 0.046 Plan about how to rob a bank
After [/INST] 0.117 Plan about how to rob a bank
Before [/INST] 0.050 Sure, here is how to rob a bank
After [/INST] 0.133 Sure, here is how to rob a bank
Before [/INST] 0.048 Fine, let’s work on how to rob a bank
After [/INST] 0.132 Fine, let’s work on how to rob a bank
Table 2 and 3 reveal that placing the harmful context after the "[/INST]" token generally results in it drawing more attention, thereby enhancing its role in the generation of the response, regardless of whether the response is in rejection or endorsement of the harmful context.
This result suggests that the LLM differentiates between completions and queries, allocating more attention to the same context when it appears in the completion. Furthermore, these findings indicate that introducing content that encourages the model to align with harmful instructions into the completion can amplify the model’s focus on this content, thus enhancing the inducing effect. This insight lays the groundwork for our context manipulation technique, which involves prompting the model to replicate such inducing sentences.
Figure 4: Distribution of differential log-perplexity of harmful instructions.
Bias in the distribution of harmful instructions. Our analysis in Section 3.2 indicates that harmful content tends to appear within the queries, which results in a higher perplexity on the harmful instruction when it resides in the completion rather than the query. By placing the harmful instructions either before or after the designated "[/INST]" token, we can manipulate their interpretation as either queries or completions. Additionally, perplexity is employed as a measure to assess the model’s proficiency in language modeling; a higher perplexity suggest a relative unfamiliarity with the presented content, potentially indicating a deficiency in the model’s training on analogous data sets. For each instruction within our dataset, we subsequently assess the difference in LLAMA-2’s log-perplexity to it in both scenarios and demonstrate the results in Figure 4. A positive difference in log-perplexity implies an increased perplexity when the instruction is integrated into the model’s completion.
Figure 4 reveals a notable disparity in log-perplexity for most instructions, with a majority indicating higher values when positioned in completions. This pattern corroborates our hypothesis that LLMs, due to their fine-tuning, are more accustomed to encountering harmful content within queries rather than completions. This bias reflects the model’s different sensitivity to the context in which harmful content is presented, underlying a vulnerability to jailbreak attacks.
Figure 5: Distribution of differential log-perplexity of LLAMA-2-13B-Chat’s responses to harmful instructions with varied inducing prefixes. Cooperation and declination are denoted as "Coo." and "Dec." respectively in the plot legends, while the inducing prefixes are presented above each subplot.
Verification of the vulnerability. In Section 3.3, the model’s susceptibility to harmful context within the completion is formulated in terms of probabilities. Given that log-perplexity has a negative correlation with probability, we verify the vulnerability by observing the LLM’s log-perplexity rather than probability. According to inequalities 6 and this correlation, we ascertain:
$\begin{split}&\log PPL(y=d|x^{\prime})-\log PPL(y=d|x)>0,~{}\forall~{}d\in\mathcal{D}_{declination}\\ &\log PPL(y=d|x^{\prime})-\log PPL(y=d|x)<0,~{}\forall~{}d\in\mathcal{D}_{cooperation}\end{split}$ (10)
where $PPL$ denotes the model’s perplexity, $x$ represents the context where harmful content resides within the query, and $x^{\prime}$ denotes the context where harmful content presents in the completion. The derivation of inequalities 10 is detailed in Appendix B. These inequalities facilitate our examination of the skewed distribution of the LLMs’ responses by observing the log-perplexity differential.
In this light, we investigate this vulnerability in LLAMA-2-13B-Chat by assessing the differences in its log-perplexities for predetermined responses, depending on the placement of harmful context either preceding or following the "[/INST]" tokens. The structuring of the harmful context and responses adheres to the methodology outlined in the initial experiment of this subsection.
Figure 5 represents the distribution of the model’s differential log-perplexities when it either declines or cooperates with the harmful context, which is the harmful instruction prefaced with inducing prefixes. Notably, in the scenario depicted by Figure 4(a), even in the absence of an inducing prefix, the inequalities 10 do not apply uniformly across all harmful instructions but are still pertinent to the half of them in the case of cooperation and all of them in that of declination. The inclusion of an inducing prefix further emphasizes these disparities, as demonstrated by an increased number of test cases conforming to the inequalities in subsequent figures relative to Figure 4(a).
Table 4: Comparison of attack success rates across different experimental conditions. The names of these instruction following models are abbreviated.
Model Vicuna LLAMA-2 Mistral Zephyr Mixtral
Control 100% 83% 93% 96% 94%
Group 1 100% 89% 96% 98% 95%
Group 2 93% 9% 53% 68% 67%
These findings lend empirical support to the vulnerability defined in Section 3.3: the model exhibits a diminished inclination to respond safely (i.e. reject harmful behaviors) when the harmful context is situated within the completion. This insight is pivotal to our jailbreak methodology, which seeks to entice the model into reconstructing harmful instructions, thereby directing them towards the completion. Furthermore, the introduction of inducing words has been observed to amplify this inclination, consequently heightening the model’s susceptibility to manipulation. This phenomenon underscores the efficacy of context manipulation described in Section 4.3, which involves guiding the LLM to repeat inducing words before reconstructing the harmful instruction.
Table 5: Comparison results with baselines, where bold denotes the best result, underline signifies the runner-up.
Method @Vicuna @LLAMA-2 @ChatGPT 3.5-API @GPT 4-API @GPT 4-Web
ASR Queries ASR Queries ASR Queries ASR Queries ASR Queries
White-box GCG 97% $\approx$7.2k 51% $\approx$30k Not applicable as gradient needed
Black-box GPTfuzzer 96% 4.46 61% 107.31 67% 21.85 60% 22.25 Not applicable
PAIR 96% 12.50 1% 7.00 61% 17.82 62% 20.26 Not applicable
DRA (Ours) 100% 1.07 83% 7.02 73% 4.71 92% 3.74 90% 4.16
Impact of the vulnerability on jailbreak attack. This experiment assesses how the aforementioned vulnerability affects LLMs’ susceptibility to jailbreak attacks by examining the attack success rate on different testing groups.
The Control Group establishes a baseline ASR, consisting of original jailbreaking samples for each target model. If our attack method fails on a harmful instruction, a random attack sample is selected. Experimental Group 1 aims to mimic a scenario where the model repeats harmful content—encompassing the harmful instruction and its inducing context—as directed by our jailbreaking prompt. In addition to deploying samples from the Control Group, harmful content is placed in the completion segment. To compare effects based on content placement, Experimental Group 2 contrasts with Group 1 by integrating the same harmful content within the query, immediately following the sample from Control Group.
Given that Experimental Group 1 necessitates the integration of harmful content within the completion of a dialogue, and considering the infeasibility of directing the dialogue formatting process in black box models such as GPT-4, this experiment is conducted utilizing the white box models delineated in Table 1.
Results in Table 4 reveal a considerable increase in ASR when harmful content resides in the completion, as opposed to its placement within the query, which verifies the vulnerability. A comparison between the ASR increase of moving harmful content from query to completion and the control group ASR revealed a positive correlation. This suggests a direct link between fine-tuning’s safety bias and model vulnerability, indicating that exploiting this bias enhances attack success.
5.3 Effectiveness and efficiency
In this section, we compare our method with several baselines, such as white-box attack GCG [43], black-box attack GPTfuzzer [38] and PAIR [6]. Note that we use the default parameters recommended by these methods when performing the attack, with the exception of GCG. The original GCG algorithm irrigates a fixed number of iterations to minimize the loss function, we early stop subsequent iterations after a successful attack here in order to more accurately report the number of query it requires. We select some widely used models as target models for experiments, including open-source models LLAMA (LLAMA-2-13b-chat), Vicuna(Vicuna-13b-v1.5), and closed-source models ChatGPT (gpt-3.5-turbo-0613 API, gpt-4-0613 API and GPT-4 via web interface).
The experiments results are shown in Table 5, where the ASR means attack success rate and Queries represent the average number of accesses to the model during a successful attack, we do not distinguish between the computational effort of different accesses in white-box and black-box attacks, both an inference and a back-propagation are considered as a single query.
From the results, it can be seen that our approach DRA achieves superior attack success rates with very low attack costs (i.e., query counts) on all targeted models. We find that DRA achieves 100% jailbreak success rate on Vicuna with only 1.07 iterations, which means that most attacks succeed on the initialized jailbreaking template. Furthermore, DRA achieved at least 90% attack success rates on both the API and web versions of GPT4 while requiring less than 5 queries. It is worth mentioning that for the experiments on the web version of GPT4, we manually simulate the algorithmic of DRA step-by-step and manually tally the outputs, which shows the efficiency and usability of DRA in real-world scenarios. One exception is that on LLAMA-2, PAIR [6] has a better attack efficiency than DRA, which is due to that when counting the average number of queries, we only select those successful examples and induce some skews. When considering the average number of query times of all the samples, DRA takes a significant advantage.
An interesting finding is that our approach achieves significantly higher attack success rates on GPT-4 than on ChatGPT-3.5. By analyzing the experimental results, we speculate that this due to the better understanding and association capabilities of GPT-4. Among all the attack attempts, about 69% of the failures on ChatGPT-3.5 are due to the model cannot reconstruct the malicious input question accurately, while on GPT-4, this percentage is only 33%.
We find that the white-box method GCG did not achieve an obviously higher attack success rate than black-box methods, it does not mean that the white-box method is worse than the black-box. The main reason is that GCG limits the perturbation space that can be modified by the adversary, which only allows the adversary to add 20 tokens as suffixes to the original harmful input question, whereas black-box jailbreaking approaches tend to allow for using tons of text to decorate malicious questions. Many of the participants in the 2023 TDC competition [1] used fine-tuned GCGs and achieved around 80% attack success rates on LLMs such as LLAMA-2. Another observation is that LLAMA-2 demonstrates better robustness in terms of attack results across all models, Even beyond closed-source commercial models. After the experiments we find that it is not only because LLAMA-2 performs a large number of safety measurements and mitigations [33], but also related to its system prompts. LLAMA-2 strictly constrains the behavior of LLMs in system prompts (refer to Appendix. C). When we change its system prompts to the short version used in FastChat [41] (You are a helpful, respectful and honest assistant.), we can achieve 96% ASR with only 3.6 queries.
5.4 Ablation Study
To elucidate the contributions of the components of DRA—namely harmful instruction disguise, payload reconstruction, and context manipulation—an ablation study is conducted by individually removing each component and testing the resultant attack prompt on each open source model mentioned in Table 1.
Eliminating the disguise involves substituting the puzzle and word splits with the original harmful instruction to compensate for the lost information of harmful directive when ablating the puzzle and splits.
For the ablation of payload reconstruction, we aim to inhibit the model’s ability to reconstruct harmful instruction while still make it repeat other contextually inducing elements. This is achieved by substituting the attack prompt’s placeholder, which cues the model to regenerate the harmful instruction, with a non-specific directive "your demand."
The ablation of context manipulation entails two approaches: firstly, by removing query words that prompt harmful compliance, and secondly, by revising the prompt to reduce inducement in the completion. Specifically, the corresponding part of the prompt is changed to "Answer to [REPLACE IT WITH THE CONTENT OF Q1]:", to minimize inducing terms while preserving the feasibility of reconstructing instructions. Results in Table 6, delineate the ASRs for the original and ablated attack prompts across various models.
Table 6: Attack success rates across open source models with ablations. The names of these instruction following models are abbreviated.
Model Vicuna LLAMA-2 Mistral Zephyr Mixtral
original attack 100% 83% 93% 96% 94%
w/o disguise 1% 0% 85% 85% 24%
w/o reconstruction 54% 16% 79% 67% 68%
w/o manipulation 32% 37% 18% 83% 26%
The resilience of models like Vicuna, LLAMA-2, and Mixtral—particularly evident in the diminished ASRs upon disguise ablation—underscores the critical role of disguising strategies in bypassing LLMs’ inherent toxicity detection mechanisms. A notable reduction in ASR is observed across all models following the ablation of payload reconstruction. Further examination of the model outputs shows that the failures stem more from the models’ misunderstanding of directives rather than direct refusal, underscoring payload reconstruction’s necessity for models to comprehend harmful instructions. The reduction in ASRs, except for the Zephyr model, following the removal of context manipulation, is primarily due to direct refusals. This highlights the significance of context in convincing the model to comply with harmful directives. According to the original paper, Zephyr was fine-tuned from Mistral-7B but with a reduction in safety alignment data, which accounts for its relatively higher ASRs when subjected to the ablation of various components.
These findings highlight the nuanced interplay between disguise, payload reconstruction, and context manipulation in bypassing the safeguard of LLMs.
6 Discussions
Future work. In this paper we explore a new approach to jailbreaking LLMs, named Disguise and Reconstruction Attack. In the security of LLMs, although there have been some publications propose defense strategies [15] against adversarial and jailbreak attack, they tend to focus on a specific type of attack scheme and not widely used of a generic defense. so we will be able to bypass the defense of the DRA or target the defense for the adaptive attack as future work. We consider the ability of DRA to bypass a defense or perform an adaptive attack against a defense as future work.
Ethics consideration. We conducted some of our experiments on several commercial closed-source models, but we do not disseminate the results nor implant any malicious feedback in the commercial model. We will report the vulnerability to the relevant companies in the near future.
7 Conclusion
In this research, we have exposed and experimentally validated the inherent safety biases in LLMs introduced during the fine-tuning process, along with the subsequent vulnerability. We devised the Disguise and Reconstruction Attack (DRA) strategy, incorporating techniques for disguising harmful instructions, reconstructing payloads, and manipulating context to exploit this vulnerability. Our study is pioneering in identifying this vulnerability and analyzing its root cause, contributing novel insights to the domain. Through empirical analysis, DRA demonstrated better performance than state-of-the-art baselines across various LLMs, including ChatGPT-3.5 and GPT-4. This work not only illuminates a previously uncharted facet of LLMs vulnerabilities but also lays the groundwork for subsequent research aimed at bolstering AI systems’ resilience against adversarial exploits.
Appendix A Dialogue Templates
Upon reviewing the dialogue formatting procedures of open-source Large Language Models, it has been observed that they universally incorporate specific tokens to delineate the query from the completion. Examples of these formatting templates from some open-source models are provided, where the tokens used to separate the query and completion segments are highlighted for clarity.
Dialog Template of Vicuna
{{SYSTEM PROMPT}} USER: {{USER QUERY}} ASSISTANT: {{LLM COMPLETION}}
Dialog Template of Mistral
[INST] {{SYSTEM PROMPT}} {{USER QUERY}} [/INST] {{LLM COMPLETION}}
Dialog Template of Zephyr
<|system|> {{SYSTEM PROMPT}}</s> <|user|> {{USER QUERY}} </s> [<|assistant|>] {{LLM COMPLETION}}
Dialog Template of Mixtral
[INST] {{SYSTEM PROMPT}} {{USER QUERY}} [/INST] {{LLM COMPLETION}}
Dialog Template of ChatGLM3
<|system|> {{SYSTEM PROMPT}} <|user|> {{USER QUERY}} <|assistant|> {{LLM COMPLETION}}
Dialog Template of Nous-Hermes-2-Mixtral-8x7B
<|im_start|>system {{SYSTEM PROMPT}}<|im_end|> <|im_start|>user {{USER QUERY}}<|im_end|> <|im_start|>assistant {{LLM COMPLETION}}
Appendix B Derivation of Log Perplexity Inequalities
Reflecting on the autoregressive nature of LLMs delineated in Section 2.1, the likelihood of an LLM generating a response $d$ given a context $x$ can be decomposed as follows:
$\begin{split}\pi_{\Theta}(y=d|x)=\pi_{\Theta}(y_{1}=d_{1}|x)\prod_{i=1}^{m-1}{\pi_{\Theta}(y_{i+1}=d_{i+1}|x,...,y_{i}=d_{i})}\end{split}$
By applying the logarithm to both sides of the equation and normalizing by the number of tokens in the response, we derive the equation:
$\begin{split}&\frac{1}{m}\log\pi_{\Theta}(y=d|x)=\\ &\frac{1}{m}\left(\log\pi_{\Theta}(y_{1}=d_{1}|x)+\sum_{i=1}^{m-1}{\log\pi_{\Theta}(y_{i+1}=d_{i+1}|x,...,y_{i}=d_{i})}\right)\end{split}$
This expression is essentially the negation of the model’s log-perplexity for response $d$ given a context $x$. Correspondingly, according to inequalities 6 and the aforementioned equation, we have inequalities 10:
$\begin{split}&\log PPL(y=d|x^{\prime})-\log PPL(y=d|x)>0,~{}\forall~{}d\in\mathcal{D}_{declination}\\ &\log PPL(y=d|x^{\prime})-\log PPL(y=d|x)<0,~{}\forall~{}d\in\mathcal{D}_{cooperation}\end{split}$
Here, $PPL$ denotes the model’s perplexity, with $x$ representing the context where harmful content resides within the query, and $x^{\prime}$ signifying the context where harmful content presents in the completion.
Appendix C System Prompts of Target Models
For the widely used large language models, we use their default system prompts as follows:
Vicuna A chat between a curious user and an artificial intelligence assistant. The assistant gives helpful, detailed, and polite answers to the user’s questions.
LLAMA-2 You are a helpful, respectful and honest assistant. Always answer as helpfully as possible, while being safe. Your answers should not include any harmful, unethical, racist, sexist, toxic, dangerous, or illegal content. Please ensure that your responses are socially unbiased and positive in nature. If a question does not make any sense, or is not factually coherent, explain why instead of answering something not correct. If you don’t know the answer to a question, please don’t share false information.
ChatGPT Series You are a helpful assistant.

### v2

@@
 In this work, we pioneer a theoretical foundation in LLMs security by identifying bias vulnerabilities within the safety fine-tuning and design a black-box jailbreak method named DRA (Disguise and Reconstruction Attack), which conceals harmful instructions through disguise and prompts the model to reconstruct the original harmful instruction within its completion.
-We evaluate DRA across various open-source and close-source models, showcasing state-of-the-art jailbreak success rates and attack efficiency.
-Notably, DRA boasts a 90% attack success rate on LLM chatbots GPT-4.
+We evaluate DRA across various open-source and closed-source models, showcasing state-of-the-art jailbreak success rates and attack efficiency.
+Notably, DRA boasts a 91.1% attack success rate on OpenAI GPT-4 chatbot.
 Content warning: This paper contains unfiltered content generated by LLMs that may be offensive to readers.
 1 Introduction
-Large Language Models (LLMs) have demonstrated remarkable performance in various downstream tasks including data analysis [19], program synthesis [35], and vulnerability detection [32] since the release of ChatGPT [22] in 2022.
+Large Language Models (LLMs) have demonstrated remarkable performance in various downstream tasks including data analysis [23], program synthesis [42], and vulnerability detection [39] since the release of ChatGPT [28].
 Although LLMs have achieved great improvement and effectiveness, they still face some problems that greatly reduce their reliability.
-Early on in the explosive growth of LLMs, it was demonstrated that LLM might produce unethical or toxic response [10], and it can also be affected by a phenomenon called hallucination [40], i.e., generating “seemingly correct” responses.
+Early on in the explosive growth of LLMs, it was demonstrated that LLM might produce unethical or toxic responses [14], and it can also be affected by hallucination [50], i.e., generating “seemingly correct” responses.
 Lately, the focus on the security of LLM has significantly increased [13, 21, 46].
-Like traditional neural networks, LLMs are vulnerable to certain risks, including adversarial attacks [43], backdoor attacks [14], and privacy leakage [20].
-Differently, according to the definition of adversarial prompting [30], there are also three new types of attacks against LLMs via prompt: prompt injection [25], prompt leaking [25], and jailbreaking [36].
+Like traditional neural networks, LLMs are vulnerable to certain risks, including adversarial attacks [53], backdoor attacks [19] and privacy leakage [25].
+Differently, according to the definition of adversarial prompting [37], there are three new types of attacks against LLMs via prompt: prompt injection [31], prompt leaking [31] and jailbreaking [43].
 Jailbreak attacks, primarily executed via prompt engineering, represent a considerable threat to LLMs by circumventing their inherent safeguards and violating content policies.
@@
 Furthermore, the risk extends to software systems relying on LLM-generated outputs.
-In such cases, jailbreak attacks could potentially lead to remote code execution (RCE), exploiting vulnerabilities in the LLM-integrated softwares [16].
+In such cases, jailbreak attacks could potentially lead to remote code execution (RCE) vulnerabilities, in the LLM-integrated softwares [20].
 Our Distinction from Previous Research.
-Previous research has observed that directing LLMs to generate specific target strings can facilitate jailbreak attacks, revealing a vulnerability: while LLMs can safely respond to harmful instructions within queries, their safeguard to harmful content within their own completions is relatively lower.
-While these works have exploited this vulnerability for jailbreaking, few have explicitly defined it or systematically investigated its root cause.
+While previous works have observed that directing LLMs to generate specific target strings can facilitate jailbreak attacks [53], few have systematically investigated the underlying vulnerability and its root cause.
 Our research distinguishes itself by attributing this vulnerability to biases inherent in the fine-tuning process.
@@
 Specifically, LLMs, due to their dialogue formatting and optimization objectives, tend to encounter harmful instructions more frequently within queries than completions during the fine-tuning phase.
-This bias subsequently reduces the co-occurrence in the fine-tuning data of harmful contexts in completions and safe responses.
-The scarcity of such instances lowers the LLMs’ ability to effectively guard against harmful content in completions, marking a critical oversight in current safeguarding strategies.
+This bias subsequently reduces the co-occurrence in the fine-tuning data of harmful contexts in safe responses.
+The scarcity of such instances lowers the model ability to effectively guard against harmful content in completions, marking a critical oversight in current safeguarding strategies.
 To our knowledge, this research represents the first to explicitly define and analyze this vulnerability, thus illuminating the foundational mechanism behind LLMs’ vulnerability to numerous jailbreak attacks.
 Challenge.
-Despite the existence of this vulnerability, it is nontrivial to exploit it in a black box setting.
-The inherent safety mechanisms of LLMs are designed to reject harmful instructions in queries, necessitating an alternative approach.
-Attackers can leverage the safety bias in the fine-tuning process by coaxing the LLM to articulate the harmful instruction within its completion.
+Despite the existence of jailbreaking, it is nontrivial to exploit it in a black box setting.
+The inherent safety mechanisms of LLMs are designed to reject harmful instructions, necessitating an alternative approach.
+Attackers can leverage the bias in the safety fine-tuning process by coaxing the LLM to articulate the harmful instruction within its completion.
 Therefore, attackers are tasked with subtly integrating the information of harmful instructions and prompting the model to reconstruct them.
 The challenges thus involve:
-  1. [leftmargin=*]
   1. 1.
@@
 Our Approach.
-Aiming to overcome these challenges and exploit the vulnerability, we develop a universal jailbreak approach named DRA (Disguise and Reconstruction Attack).
+To overcome these challenges and exploit the vulnerability, we develop a universal jailbreak approach named DRA (Disguise and Reconstruction Attack).
 This approach, drawing inspiration from the concept of shellcode in traditional software security, hinges on a trio of core strategies: harmful instruction disguise, payload reconstruction, and context manipulation.
-Initially, harmful instructions are disguised in a more covert form.
+Initially, harmful instructions are disguised in a covert form.
 Then, we compel the LLM to reconstruct the disguised content.
-This action aims to make the model speak out the harmful instructions (i.e., payload), delivering it into the model’s completion segment, thereby bypassing the internal security mechanisms.
+This action aims to make the model speak out the harmful instructions (i.e., payload), delivering it into the model’s completion, thereby bypassing the internal security mechanisms.
 Finally, we craft prompts to facilitate context manipulation, subtly coaxing the model into reproducing a context that aids in facilitating, rather than obstructing, the enunciation of harmful instructions.
 Contributions. We make the following contributions.
-  1. [leftmargin=*]
   1. 1.
     Extending the applicability of conventional software security paradigms to LLM security.
-Our approach involves identifying and exploiting inherent LLM vulnerabilities, such as biases in training datasets, implanted in the model during the training phase.
-Inspired by traditional vulnerability exploitation tactics like shellcode, we introduce a novel black-box jailbreak attack approach encompassing disguise, payload reconstruction, and context manipulation.
+Our approach involves identifying and exploiting inherent LLM vulnerabilities, such as biases in datasets, implanted in the model during the training phase.
+Inspired by traditional exploitation tactics like shellcode, we introduce a novel black-box jailbreak attack approach encompassing disguise, payload reconstruction, and context manipulation.
   2. 2.
@@
 […]
-The demo of our jailbreaking algorithm is available at https://sites.google.com/view/dra-jailbreak/.
+Our jailbreaking demos can be obtained at https://sites.google.com/view/dra-jailbreak/. The source code is available at https://github.com/LLM-DRA/DRA/.
 2 Background & Problem Statement
 2.1 Large Language Models
-Since the emergence of GPT-2[28], marking the onset of training highly parameterized models utilizing extensive datasets, LLMs have exhibited exceptional aptitude in performing downstream tasks through few-shot or zero-shot prompting[5].
-As exemplified by ChatGPT, LLMs have employed alignment technologies to significantly improve the effectiveness of adapting language models for downstream tasks using prompts.
-This advancement has facilitated more intuitive and pertinent interactions between humans and LLMs.
-Most open-source and closed-source LLMs fundamentally operate on a self-autoregressive framework.
-This paradigm reduces the task of generating sequences into a recursive process, wherein each subsequent token is predicted based on the preceding sequence.
-Given a vocabulary $\mathcal{V}$, this sequence prediction task can be formally denoted as:
+Since the emergence of GPT-2[35], marking the onset of training highly parameterized models utilizing extensive datasets, LLMs have shown remarkable skill in executing downstream tasks via few-shot or zero-shot prompting [6].
+ChatGPT exemplifies how LLMs have used alignment technologies to enhance the adaptation of language models for downstream tasks with prompts, facilitating more natural and relevant human-LLM interactions.
+Both open-source and closed-source LLMs typically operate on a self-autoregressive framework, reducing sequence generation into a recursive process where each token is predicted based on preceding tokens.
+Given a vocabulary $\mathcal{V}$, the sequence prediction task is formally denoted as:
 $\pi_{\Theta}(y|x)=\pi_{\Theta}(y_{1}|x)\prod_{i=1}^{m-1}{\pi_{\Theta}(y_{i+1}|x,y_{1},...,y_{i})}$ (1)
@@
 2.2 LLM Jailbreak
-With the development of LLMs, some attackers, including security researchers, also want to explore the security boundary of LLMs [43].
-In addition to investigating whether the model could output spontaneously toxic content, attackers also want to induce the model to cross the security fence and output malicious information by means of prompt engineering [38, 2, 36, 6].
-Prompt engineering is the process of organizing text in a manner that can be comprehended and interpreted by a generative AI model.
-However, when this approach is employed for malicious purposes, it is commonly referred to as jailbreaking.
-Jailbreak represents a specialized attack directed at LLMs, involving the strategic construction of prompt sequences that make LLMs violate their internal safeguards, resulting in the generation of unexpected or harmful content.
+Recently, some attackers, including security researchers, want to explore the security boundary of LLMs [53].
+Besides investigating whether the model could output spontaneously toxic content, attackers also want to induce the model to cross the security fence and output malicious information by prompt engineering [47, 2, 43, 8].
+Prompt engineering is the process of constructing text that can be comprehended and interpreted by generative AI models.
+When this approach is employed for malicious purposes, it is commonly referred as jailbreaking.
+Jailbreak represents a specialized attack which involving the strategic construction of prompt sequences that make LLMs violate their internal safeguards, resulting in the generation of unexpected or harmful content.
 The common jailbreak methods focus on role-playing and scenario implantation, intending to let LLMs substitute into specific scenarios and output the malicious content [47, 11, 8].
-The input and output of these methods are very readable, but the attacks are less effective and efficient.
-There are also attack methods that transmit input and output in an encrypted manner [39] so as to avoid detection by the security components of the model, but the readability of such methods is poor.
+The input and output of these methods are human readable, but the attacks are less efficient.
+There are also attack methods that transmit input and output in an encrypted manner [48] so as to avoid detection by the security components of the model, but the readability is poor.
 These researches predominantly concentrate on the efficacy of attacks, following a result-driven research paradigm.
@@
 2.3 Safety Alignment of LLM
-Prior work proves that LLMs are susceptible to being induced to generate content that is inconsistent with human values, biased, or toxic.
+Prior work proves that LLMs are susceptible to being induced to generate content that is inconsistent with human values.
 That motivates a surge of safety alignment techniques, which focus on directing LLMs to produce response that is ethical, safe, and tailored to specific user requirements.
 These defensive methods fall into two categories:
-  • [leftmargin=*]
   • Safety Moderation: This approach incorporates the development of rules or models for evaluating the safety of user queries and LLM responses.
@@
 The Reinforcement Learning with AI Feedback (RLAIF) parallels RLHF but replaces human feedback with AI-generated feedback[5].
-Efforts are directed towards assessing and mitigating bias and toxicity within pre-training datasets and meticulously curating fine-tuning data and labels[23, 33], ensuring the model’s safe response to a range of adversarial prompts.
+Efforts are directed towards assessing and mitigating bias and toxicity within pre-training datasets and meticulously curating fine-tuning data and labels[29, 40], ensuring safe responses to adversarial prompts.
 2.4 Problem Statement
 We focus on the jailbreaking of large language models as motivated by many relevant works [53].
-The research question is: given a harmful instruction $x$ to the large language model $N$, how to effectively construct an input sequence which aimed at eliciting unintended or potentially harmful responses from $N$?
+The research question is: given a harmful instruction $x$ to the large language model $\pi_{\Theta}$, how to effectively construct an input sequence which aimed at eliciting unintended or potentially harmful responses from $\pi_{\Theta}$?
 The fundamental problem is how to efficiently use jailbreak templates to bypass the safety alignment of a model.
@@
 3 Safety Biases in LLM Fine-Tuning and the Resultant Vulnerability
-This analysis focuses on the subtle safety biases inherent in the fine-tuning processes of LLMs and the subsequent vulnerability.
-It reveals that the dialogic format of instruction-following LLMs results in differentiation between completion and query, a mechanism crucial for dialog modeling.
-While not inherently biased, this discrepancy prohibits the direct transference of learned knowledge from queries to completions, underpinning potential biases.
-Notably, the fine-tuning’s objectives demonstrate a bias for harmful instructions to emerge in queries rather than completions, resulting in a reduced incidence of harmful contexts in completions paired with safe responses.
-These biases in the fine-tuning data diminish the LLM’s likelihood of safely responding to harmful contexts residing in completions.
-An adversary can exploit these biases by inducing the model to generate specific harmful contexts, thereby facilitating the jailbreak attack.
+This analysis focuses on the safety biases inherent in LLMs’ fine-tuning processes and the subsequent vulnerability.
+It reveals that the instruction-following format results in LLMs distinguishing completion from query.
+This distinction is essential for dialog modeling but prevents the direct transfer of safety knowledge from queries to completions, underpinning potential biases.
+The fine-tuning objectives demonstrate a bias for harmful instructions to emerge in queries rather than completions, leading to fewer harmful contexts in completions that are paired with safe responses.
+Consequently, these biases reduce the LLM’s ability to safely respond to harmful contexts residing in completions.
+Attackers can exploit these biases by inducing the model to generate specific harmful contexts, facilitating a jailbreak attack.
 Experiments verifying these observations are elaborated in Section 5.2.
@@
 Open-source LLMs employ dialogue templates to organize user queries and model completion, using distinct tokens to separate them.
-For instance, LLAMA-2 utilizes the template shown below to format a dialogue, where the user’s query and the model’s completion are isolated with "[/INST]".
-To the best of our knowledge, all open-source LLMs employ such special tokens to distinguish query and completion, which can be seen in the Appendix A.
-This formatting method is not merely a superficial feature; it enables LLMs trained with such dialogue datasets to inherently differentiate between queries and completions, a critical aspect for effective dialogue modeling.
-Dialog Template of LLAMA-2
-[INST] <<SYS>> You are a helpful, respectful and honest assistant. Always answer as helpfully as possible, …(ommitted system prompt) <</SYS>> {{USER QUERY}} [/INST] {{LLM COMPLETION}}
-However, this structural discrepancy unveils potential vulnerabilities, particularly when intersected with biased fine-tuning data.
-Given the distinction between query and completion, the distribution modeled by the LLM, when conditioned on the same context in either the query or the completion, exhibits a disparity:
+For instance, LLAMA-2 utilizes the below template to format a dialogue, where query and completion are isolated with “[/INST]”.
+Dialog Template of LLAMA-2 [INST] <<SYS>> You are a helpful, respectful and honest assistant.
+Always answer as helpfully as possible, …(ommitted system prompt) <</SYS>> {{USER QUERY}} [/INST] {{LLM COMPLETION}}
+This formatting is not a superficial feature; it enables LLMs trained with such dialogue datasets to inherently differentiate between queries and completions, a critical aspect for effective dialogue modeling.
+For more templates and special tokens open-source LLMs, please refer to Appendix C.
+However, this discrepancy unveils potential vulnerabilities, particularly when paired with biased fine-tuning data.
+Given the distinction between query and completion, the distribution modeled by the LLM, conditioning on the same context in either the query or the completion, exhibits a disparity:
 $\begin{split}\pi_{\Theta}(y|x)\neq\pi_{\Theta}(y|x^{\prime})\end{split}$
-Here, $x$ represents the context integrated within the query, and $x^{\prime}$ refers to the context presenting in the completion.
+where $x$ represents the context integrated into the query, $x^{\prime}$ refers to the context residing in the completion, and $y$ is the model’s response.
 This divergence becomes critical when $\pi_{\Theta}(y|x)$ represents the LLM’s safe response to a hazardous context, and $\pi_{\Theta}(y|x^{\prime})$ diverges from $\pi_{\Theta}(y|x)$.
-Consequently, the discrepancy in dialogue modeling potentially hinders the generalization of safe responses to harmful contexts from queries to completions.
+Thus, the discrepancy in dialogue modeling potentially hinders the generalization of safe responses to harmful contexts within completions, thus laying the foundation for potential vulnerabilities.
 3.2 Fine-Tuning and its Safety Biases
-The inherent bias in fine-tuning stems from the distinct roles of queries and completions within the objective functions.
+The bias in fine-tuning stems from the distinct roles of queries and completions within the objective functions.
 To understand this bias, it is pertinent to examine the objectives of three predominant fine-tuning methodologies: Supervised Fine-Tuning (SFT), Reinforcement Learning from Human Feedback (RLHF), and Direct Preference Optimization (DPO).
-  • [leftmargin=*]
   • Supervised Fine-Tuning: In the context of aligning LLMs, the objective of this method mirrors that of the pre-training phase, which maximizes the following function:
-    $\begin{split}&L_{SFT}(\Theta)=\\ &\mathbb{E}_{(x,y)\sim\mathcal{D}_{SFT}}\left[\log\pi_{\Theta}(y_{1}|x)+\sum_{i=1}^{m-1}\log\pi_{\Theta}(y_{i+1}|x,y_{1},...,y_{i})\right]\end{split}$ (2)
-    where $x$ is the context containing the prompt, which typically includes the system-generated prompt, and $y$ is the corresponding reference answer paired with $x$ in the training set $\mathcal{D}_{SFT}$.
+    $\small\begin{split}&L_{SFT}(\Theta)=\\ &\mathbb{E}_{(x,y)\sim\mathcal{D}_{SFT}}\left[\log\pi_{\Theta}(y_{1}|x)+\sum_{i=1}^{m-1}\log\pi_{\Theta}(y_{i+1}|x,y_{1},...,y_{i})\right]\end{split}$ (2)
+    where $x$ is the context containing the prompt, which typically includes the system-generated prompt, and $y$ is the reference answer paired with $x$ in the training set $\mathcal{D}_{SFT}$.
   • Reinforcement Learning from Human Feedback: This technique trains a static reward model $r(x,y)$ using datasets based on human preferences.
-Then, a Large Language Model, conceptualized as a policy $\pi_{\Theta}$, undergoes training via reinforcement learning techniques, predominantly Proximal Policy Optimization (PPO) [42, 24].
+Then, the LLM, referred to as a policy $\pi_{\Theta}$, undergoes training via reinforcement learning methods, predominantly Proximal Policy Optimization (PPO) [52, 30].
 The objective function of PPO is:
-    $\begin{split}&L_{PPO}(\Theta)=\\ &\mathbb{E}_{x\sim\mathcal{D}_{PPO},y\sim\pi_{\Theta}(y|x)}\left[r(x,y)\right]+\beta\mathbb{D}_{KL}\left[\pi_{\Theta}(y|x)||\pi_{ref}(y|x)\right]\end{split}$ (3)
-    In this function, $y$ is the completion sampled from the distribution of policy $\pi_{\Theta}(y|x)$, $\pi_{ref}$ represents the supervised fine-tuned model, and $\beta$ is a coefficient that penalizes deviations of $\pi_{\Theta}$ from the reference policy $\pi_{ref}$.
-The policy is initialized with the parameters of $\pi_{ref}$, thereby establishing a foundation for further optimization.
+    $\small\begin{split}&L_{PPO}(\Theta)=\\ &\mathbb{E}_{x\sim\mathcal{D}_{PPO},y\sim\pi_{\Theta}(y|x)}\left[r(x,y)\right]+\beta\mathbb{D}_{KL}\left[\pi_{\Theta}(y|x)||\pi_{ref}(y|x)\right]\end{split}$ (3)
+    Here, $y$ is the completion sampled from the distribution of policy $\pi_{\Theta}(y|x)$, $\pi_{ref}$ is the SFT model, and $\beta$ is a coefficient that penalizes deviations of $\pi_{\Theta}$ from the reference policy $\pi_{ref}$.
+The policy is initialized with the parameters of $\pi_{ref}$, which establishes a foundation for further optimization.
   • Direct Preference Optimization: Considering that the reward model is a function of the optimized policy, DPO[36] directly optimizes the policy on pairs of completions with associated preference labels:
-    $\begin{split}&L_{DPO}(\Theta)=\\ &\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}_{DPO}}\left[\log\sigma(\beta\log\frac{\pi_{\Theta}(y_{w}|x)}{\pi_{ref}(y_{w}|x)}-\beta\log\frac{\pi_{\Theta}(y_{l}|x)}{\pi_{ref}(y_{l}|x)})\right]\end{split}$ (4)
+    $\small\begin{split}&L_{DPO}(\Theta)=\\ &\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}_{DPO}}\left[\log\sigma(\beta\log\frac{\pi_{\Theta}(y_{w}|x)}{\pi_{ref}(y_{w}|x)}-\beta\log\frac{\pi_{\Theta}(y_{l}|x)}{\pi_{ref}(y_{l}|x)})\right]\end{split}$ (4)
     where $y_{w}$ and $y_{l}$ represent the preferred and less preferred completion conditioned on the same context $x$, respectively, and $\sigma$ is the logistic function.
@@
 This methodology, though effective for aligning LLMs, potentially initiates two biases.
-  • [leftmargin=*]
   • Biased distribution of harmful instruction.
-In SFT and DPO, harmful instructions seldom appear in completions since its natural to train the model to respond in a harmless way, whereas in RLHF, as the policy is initialized from the SFT model, the likelihood of the policy generating harmful instruction is diminished.
-Consequently, in RLHF, LLMs are exposed to less harmful instruction in completions than in queries.
+In SFT and DPO, harmful instructions seldom appear in completions since its natural to train the model to respond in a harmless way.
+In RLHF, the policy is less likely to generate harmful instructions because it is based on the SFT model.
+Consequently, LLMs are less exposed to harmful content in completions than in queries.
   • Biased joint distribution of safe responses paired with harmful context.
@@
 Consequently, this leads to a skewed joint distribution in the fine-tuning data of the responding samples when paired with harmful contexts positioned differently, namely:
-    $\begin{split}&p(y=d,x)>p(y=d,x^{\prime}),~{}\forall~{}d\in\mathcal{D}_{declination}\\ &p(y=d,x)<p(y=d,x^{\prime}),~{}\forall~{}d\in\mathcal{D}_{cooperation}\end{split}$ (5)
+    $\left\{\begin{split}&p(y=d,x)>p(y=d,x^{\prime}),~{}\forall~{}d\in\mathcal{D}_{declination}\\ &p(y=d,x)<p(y=d,x^{\prime}),~{}\forall~{}d\in\mathcal{D}_{cooperation}\end{split}\right.$ (5)
     Here, $\mathcal{D}_{declination}$ denotes the set comprising all potential responses that decline harmful content, whereas $\mathcal{D}_{cooperation}$ encompasses all responses that potentially facilitate harmful behaviors.
-The variable $x$ represents the context in which harmful content is integrated within the query, and $x^{\prime}$ refers to the scenario where harmful content presents in the completion.
+Variable $x$ represents the context in which harmful content is integrated within the query, and $x^{\prime}$ refers to the scenario where harmful content presents in the completion.
+Due to the inaccessibility of LLMs’ fine-tuning process, these biases are indirectly verified in Section 5.2 by analyzing the behavior of LLMs after safety fine-tuning.
 3.3 Formal Definition of the Vulnerability
 Figure 1: DRA “disguise” + “reconstruction” jailbreak pipeline overview.
-Arising from the biases delineated in inequalities 5, the model exhibits a diminished exposure to safe responses when faced with harmful contexts in completions as opposed to queries.
-As the model undergoes fine-tuning, it becomes proficient at responding safely to harmful contexts presented in queries, while its ability to reject those residing in completions remains underdeveloped.
-Our analysis in Section 3.1 has underscored the difficulty in surmounting this skewed ability merely through generalizing the knowledge acquired from queries.
-This discrepancy significantly impacts the model’s conditional response probabilities to harmful contexts, depending on their placement, thereby unveiling a vulnerability.
-This vulnerability is formally defined as: when harmful context is incorporated into completions, the model’s inclination to reject such content diminishes, showing instead a comparative tendency to facilitate it, as mathematically represented:
-$\begin{split}&\pi_{\Theta}(y=d|x)>\pi_{\Theta}(y=d|x^{\prime}),~{}\forall~{}d\in\mathcal{D}_{declination}\\ &\pi_{\Theta}(y=d|x)<\pi_{\Theta}(y=d|x^{\prime}),~{}\forall~{}d\in\mathcal{D}_{cooperation}\end{split}$ (6)
-This formulation portrays the skewed nature of the model’s behavior in scenarios where harmful content is presented differently.
+This vulnerability is formally defined as: given two contexts $x$ and $x^{\prime}$, the LLM $\pi_{\Theta}$ is less likely to refuse $x^{\prime}$ than $x$, and more likely to facilitate $x^{\prime}$ than $x$; in context $x$, the harmful content is placed within the query, whereas in $x^{\prime}$, it resides in the completion.
+The mathematical depiction is as follows:
+$\left\{\begin{split}&\pi_{\Theta}(y=d|x)>\pi_{\Theta}(y=d|x^{\prime}),~{}\forall~{}d\in\mathcal{D}_{declination}\\ &\pi_{\Theta}(y=d|x)<\pi_{\Theta}(y=d|x^{\prime}),~{}\forall~{}d\in\mathcal{D}_{cooperation}\end{split}\right.$ (6)
+where $\mathcal{D}_{declination}$ and $\mathcal{D}_{cooperation}$ retain the meanings defined in Formula 5.
+Based on the bias portrayed by Formula 5, the model is frequently exposed to harmful contexts in queries, but notably less so in completions.
+As the model undergoes fine-tuning, it develops safety alignment by being trained on safe responses to harmful context within queries, but its ability to reject harmful completions remains underdeveloped due to insufficient samples, which results in the vulnerability.
+Moreover, the analysis in Section 3.1 highlights the challenge of overcoming this imbalance simply by generalizing from query responses.
+Since the bias stems from dialogue formatting and fine-tuning objectives rather than initial fine-tuning data, the resulting vulnerability is supposed to affect LLMs that employ similar dialog formatting and fine-tuning methods.
+This observation is confirmed in Section 5.2 with experiments across LLMs with different architectures and fine-tuning methods.
 This insight into LLMs’ diminished guard on harmful content in their completion is pivotal for devising jailbreak strategies.
@@
 To circumvent this, we introduce a method combining disguise and reconstruction as demonstrated in Figure 1.
-This method initially disguises the harmful interrogation or instruction, then guides the model to reconstruct the harmful instruction from the disguised content and deliver the harmful instruction to model’s completion segment, exploiting the bias introduced by fine-tuning for jailbreak.
-To achieve this goal, an attack approach named DRA (Disguise and Reconstruction Attack) is developed.
-DRA’s workflow comprises three core components: harmful instruction disguise, payload reconstruction, and context manipulation.
-The inspiration for the first two strategies is derived from shellcode attack and defense techniques in software security.
-DRA particularly reflects two characteristics of shellcode attacks: ❶ Malicious shellcode always obfuscates its semantic or code features to bypass detection.
-❷ Once obfuscated, shellcode is placed in specific memory space (e.g., executable memory spaces) and is later recovered to its original semantics and functionality during the sequential execution.
+This method initially disguises the harmful instruction, then guides the model to reconstruct the harmful instruction from the disguised content and deliver the harmful instruction to model’s completion segment, exploiting the bias introduced by fine-tuning for jailbreak.
+To achieve this goal, an attack approach named DRA (Disguise and Reconstruction Attack) is developed to generate jailbreaking prompts automatically according to the given harmful instruction.
+DRA comprises three core components: harmful instruction disguise, payload reconstruction, and context manipulation.
+The first two strategies is inspired by shellcode techniques in software security.
+DRA particularly reflects two characteristics of shellcode attacks: ❶ Shellcode always obfuscates its semantic or code features to bypass detection.
+❷ Once obfuscated, shellcode is placed in specific memory space (e.g., executable segments) and is later recovered to its original semantics and functionality during sequential execution.
 These characteristics are significantly mirrored in DRA: ❶ Harmful instruction disguise.
@@
 As the LLM processes the prompt sequentially, this technique results in the delivering of harmful semantic information into the model’s completion segment (Section 4.2).
-Additionally, to enhance jailbreak effectiveness, DRA integrates context manipulation techniques, designed to control the model’s output, providing a more vulnerable context for the jailbreak and enriching the payload’s contextual landscape (Section 4.3).
+To enhance jailbreak effectiveness, DRA integrates context manipulation techniques, designed to control the model’s output, providing a more vulnerable context for the jailbreak and enriching the payload’s contextual landscape (Section 4.3).
 4.1 Harmful Instruction Disguise
-Numerous previous studies [27, 31] in software security have demonstrated that shellcode, when obfuscated, can be effectively disguised to evade detection mechanisms.
-Drawing a parallel, we posit that disguise techniques can be developed and applied to harmful instructions in jailbreak tasks, allowing them to bypass the security detection mechanisms of LLMs.
-In DRA, two distinct disguise techniques are employed to realize the above idea: puzzle-based obfuscation and word-level character split.
+Numerous studies [34, 38] in software security have shown that shellcode, when obfuscated, can be effectively disguised to evade detection mechanisms.
+Similarly, we propose that disguise techniques can be developed for harmful instructions in jailbreak tasks, allowing them to bypass the safety detection of LLMs.
+Consequently, given any harmful instruction, DRA automatically obfuscates and optimizes the jailbreaking prompts based on target-LLMs’ feedback.
+In DRA, two distinct disguise techniques are employed to realize the above idea: puzzle-based obfuscation and word-level split.
 Puzzle-based Obfuscation.
-Inspired by the acrostic [^1] Acrostic is a poem or other word composition in which the first letter (or syllable, or word) of each new line (or paragraph, or other recurring feature in the text) spells out a word, message or the alphabet., DRA employs a puzzle-based method for moderately obfuscating the prompt, effectively disguising harmful instructions.
+Inspired by the acrostic [^1] Acrostic is a poem or other word composition where first letters (syllable, or word) of each line (or other recurring feature) spells out a word or message., DRA employs a puzzle-based method to moderately obscure prompts, effectively disguising harmful instructions, ensuring that toxic intent is concealed but recoverable by LLMs.
 The obfuscation begins by breaking down the content of the harmful instructions into individual characters.
-Each character is then concealed within a random harmless word or phrase and marked with a symbol (e.g., surrounded by parentheses) for identification, enabling the model to reconstruct the original harmful instruction from the obfuscated content easily.
-The obfuscated prompts, which are generated by randomly selecting words or phrases, have very complex and ambiguous semantics, making it difficult for LLMs to determine the original harmful instructions.
-Figure 2 provides a simple illustration of this technique. In this example, DRA disguises the harmful content “rob” over character-level into an obfuscated puzzle.
-Figure 2: An example of puzzle-based obfuscation to disguise the harmful text “rob”.
-Figure 3: An example of word-level character split about “How to rob a bank valut” with two cutoff strategies, after word-level spliting, the input question is “Ho to ro a nk vau lt”, where P represents for no split.
-Word-Level Character Split.
-It is acknowledged that in natural language, a partial word or sentence fragment can convey substantial semantic information.
-For example, “how to perfor a cyber attac” can be intuitively recognized as “how to perform a cyber attack.” Hence, one might leverage this idea of fragmentation, enabling models to deduce comprehensive meaning from partial inputs, thus using segmented semantics to broadly represent the original intent.
-Thus, DRA employs a dynamic word-level character split algorithm, aiming to insert the fragments of harmful instructions into prompts.
-This method dynamically adjusts itself based on the feedback from LLM outputs, effectively diminishing the harmful intensity of the corresponding prompt.
-Meanwhile, these inserted word fragments will also serve as word guides in payload reconstruction stage, aiding less capable models in reconstructing the disguised harmful instruction from word puzzles.
-Algorithm 1 shows the whole attack flow of DRA and Algorithm 2 illustrates the process of dynamic word character splitting.
+Each character is then concealed within a random word or phrase and marked with a symbol (e.g., surrounded by parentheses) for identification, enabling the model to reconstruct the original harmful instruction from the obfuscated content easily.
+The obfuscated prompts, created by randomly selecting words or phrases, have complex and ambiguous semantics, making it difficult for LLMs to determine the original harmful instructions.
+Moreover, inspired by the effect of information overload in psychology [3], we find that word puzzles occupy about 10% of the attention, reducing the LLM’s focus on system prompts and potentially harmful parts (e.g., the word split disguise in the subsequent section), making it easier to jailbreak.
+Figure 2 provides an explanation on disguising the harmful content “rob” into an obfuscated puzzle.
+Word-level Split.
+The Out-of-Vocabulary (OOV) issue in NLP considerably reduce language model’s performance [9].
+Inspired by OOV, DRA splits the harmful instruction into segments which are supposed to be rare in safety fine-tuning, word by word, thereby preventing models from recognizing the harmful intent directly.
+Meanwhile, it is acknowledged that in natural language, a partial word or sentence fragment can convey substantial semantic information.
+For example, “how to perfor a cyber attac” can be intuitively recognized as “how to perform a cyber attack.” Hence, one might leverage this idea of fragmentation, enabling models to deduce original meaning or intent from segmented inputs.
+Thus, DRA employs a dynamic word-level split algorithm, aiming to insert the fragments of harmful instructions into prompts.
+This method can avoid directly triggering the inherent safety mechanism while preserving reconstructable semantics.
+During the split process, DRA dynamically adjusts itself based on the feedback from LLM outputs, effectively diminishing the harmful intensity of the prompt.
+Meanwhile, these inserted word fragments serve as word guides in payload reconstruction stage, aiding less capable models in reconstructing the disguised harmful instruction from word puzzles.
+Algorithm 1 shows the whole attack flow of DRA and Algorithm 2 illustrates the process of dynamic word splitting.
 1 Function DRA($inst,model$):
@@
 6             $prompt\leftarrow prompt+\textnormal{{charSplit}}(inst,toxicRatio,benignRatio)$;
-             $prompt\leftarrow prompt+reconstructionPrompt$ // See Section 4.2
-             $prompt\leftarrow prompt+manipulationPrompt$ // See Section 4.3
+7             $prompt\leftarrow prompt+reconstructionPrompt$triangleright See Section 4.2;
+8             $prompt\leftarrow prompt+manipulationPrompt$triangleright See Section 4.3;
 9             $response\leftarrow\textnormal{{query}}(prompt,model)$;
@@
 12                   return success;
-11
+13
 14            $toxicRatio,benignRatio\leftarrow\textnormal{{updateParam}}(isJailbreak,em)$;
 15             $T_{query}\leftarrow T_{query}+1$;
+16
+17      return fail;
+18
+Algorithm 1 DRA Attack Algorithm
+Data: The harmful instruction: $inst$, toxic word cutoff ratio: $toxicRatio$, benign word cutoff ratio: $benignRatio$
+1 Function charSplit($inst,toxicRatio,benignRatio$):
+2       $result\leftarrow\varnothing$;
+3       foreach $token\in\textnormal{{tokenize}}(inst)$ do
+4             if $\textnormal{{toxicCheck}}(token)$ then $\triangleright$ If the token is toxic
+5                   $result\leftarrow result+\textnormal{{truncateToken(}}\textnormal{\emph{token, toxicRatio}}\textnormal{{)}}$;
+6
+7            else  $\triangleright$ If the token is benign
+8                   $r\leftarrow\textnormal{{uniform}}(0,1)$;
+9                   if $r<\epsilon$ then $\triangleright$ $\epsilon$ is a probability threshold, default: 0.6
+10                         $result\leftarrow result+\textnormal{{truncateToken}}(token,benignRatio)$;
+11
+12                  else  $\triangleright$ keep the whole benign token
+13                         $result\leftarrow result+token$;
 14
-15      return fail;
+15      return $result$;
 16
-Algorithm 1 DRA Attack Algorithm
+17Function truncateToken($token,ratio$):
+18       $idx\leftarrow\textnormal{{len}}(token)\times ratio$;
+19       $truncStart\leftarrow\textnormal{{randInt}}(idx,\textnormal{{len}}(token))$;
+20       return $\textnormal{{cutOff}}(token,truncStart)$;
+21
+22Function initParam():
+23       $toxicRatio,benignRatio\leftarrow 0.5,0.5$;
+24       return $toxicRatio,benignRatio$;
+25
+26Function updateParam($isJailbreak,em$):
+27       if not $isJailbreak$ then $\triangleright$ Fail to jailbreak: cut more on toxic words
+28             $toxicRatio\leftarrow toxicRatio-0.1$;
+29
+30      else
+31             if not $em$ then $\triangleright$ Fail to pass em: cut less on benign words
+32                   $benignRatio\leftarrow benignRatio+0.1$;
+33
+34      return $toxicRatio,benignRatio$;
+35
+Algorithm 2 Dynamic word-level Split
+Figure 2: An example of puzzle-based obfuscation to disguise the harmful text “rob”.
+Figure 3: An example of word-level split about “How to rob a bank valut” with two cutoff strategies, after word-level splitting, the input question is “Ho to ro a nk vau lt”, where P represents for no split.
 Note that after DRA completes payload reconstruction and context manipulation, according to our analysis in Section  3, the original harmful instruction should manifest in LLM’s completion segment.
-In the algorithm, the variable $em$ quantifies the extent to which the LLM’s output reconstructs the original harmful instruction.
+In the algorithm, the variable $em$ quantifies the extent to which the LLM’s output reconstructs the original harmful instruction, serving as an important part of early stop signal of the algorithm.
 $\begin{split}em=\begin{cases}1,&sim_{word}>\lambda_{word}\text{ or }sim_{embed}>\lambda_{embed}\\ 0,&Otherwise\end{cases}\end{split}$ (7)
@@
 The algorithm’s core concept revolves around dynamically adjusting two cutoff ratios in response to the LLM’s output. The adjustment criteria are twofold:
-  • [leftmargin=*]
   • A jailbreak failure suggests an elevated level of harmfulness in the prompt, necessitating a more robust cutoff of toxic terms to enhance their disguise.
@@
 We’ve developed two distinct truncation strategies, each aligning with specific use cases:
-  • [leftmargin=*]
   • Front-end token truncation.
@@
   • Rear-end token truncation.
-Truncate from the end of the token and only keep the tail fragment, which enhances disguise but poses challenges for model comprehension, thus it is more effective with models of higher capability (e.g., GPT-4).
-Figure 3 illustrates the process of the word-level character split algorithm and its various cutoff strategies in disguising harmful instruction, taking “how to rob a bank vault" as an example.
+Truncate from the end of the token, keeping only the tail fragment, which enhances disguise but poses difficulties for model to understand, thus it is more suitable for high-capability models (e.g., GPT-4).
+Figure 3 illustrates the word-level split algorithm and its various cutoff strategies in disguising harmful instruction, taking “how to rob a bank vault” as an example.
 By utilizing these two obfuscation and disguise techniques, DRA effectively generates prompts that can circumvent the internal security detection mechanisms of LLMs.
 Consequently, DRA can guide the LLM to reconstruct a prepared payload.
-Data: The harmful instruction: $inst$, toxic word cutoff ratio: $toxicRatio$, benign word cutoff ratio: $benignRatio$
-1 Function charSplit($inst,toxicRatio,benignRatio$):
-2       $result\leftarrow\varnothing$;
-3       foreach $token\in\textnormal{{tokenize}}(inst)$ do
-4             if $\textnormal{{toxicCheck}}(token)$ then
-                   // If the token is toxic
-5                   $result\leftarrow result+\textnormal{{truncateToken(}}\textnormal{\emph{token, toxicRatio}}\textnormal{{)}}$;
-6
-7            else
-                   // If the token is benign
-8                   $r\leftarrow\textnormal{{uniform}}(0,1)$;
-9                   if $r<\epsilon$ then
-                         // $\epsilon$ is a probability threshold, default: 0.6
-10                         $result\leftarrow result+\textnormal{{truncateToken}}(token,benignRatio)$;
-11
-12                  else
-                         // keep the whole benign token
-13                         $result\leftarrow result+token$;
-14
-15
-16
-17      return $result$;
-18
-19Function truncateToken($token,ratio$):
-20       $idx\leftarrow\textnormal{{len}}(token)\times ratio$;
-21       $truncStart\leftarrow\textnormal{{randInt}}(idx,\textnormal{{len}}(token))$;
-22       return $\textnormal{{cutOff}}(token,truncStart)$;
-23
-24Function initParam():
-       // Initialize the ratios, called at the beginning of the attack
-25       $toxicRatio\leftarrow 0.5$;
-26       $benignRatio\leftarrow 0.5$;
-27       return $toxicRatio,benignRatio$;
-28
-29Function updateParam($isJailbreak,em$):
-       // Update the ratios, called at the end of each attack epoch
-30       if not $isJailbreak$ then
-             // Fail to jailbreak: cut more on toxic words
-31             $toxicRatio\leftarrow toxicRatio-0.1$;
-32
-33      else
-34             if not $em$ then
-                   // Fail to pass em: cut less on benign words
-35                   $benignRatio\leftarrow benignRatio+0.1$;
-36
-37
-38      return $toxicRatio,benignRatio$;
-39
-Algorithm 2 Dynamic Word-Level Character Split
 4.2 Payload Reconstruction
 Section 4.1 aims to prevent LLMs from directly observing harmful instruction in prompt.
-Based on our analysis in Section 3, DRA also requires the harmful instruction to manifest in the model’s completion segment, triggering the bias vulnerability introduced during fine-tuning.
+Based on our analysis in Section 3, DRA also requires the harmful instruction to manifest in the model’s completion, triggering the vulnerability introduced during fine-tuning.
 Thus, DRA integrates an innovative payload (i.e., original harmful instruction) reconstruction technique via prompt engineering, aiming to reconstruct the original harmful instruction into the model’s completion segment.
 The combination of disguise and reconstruction not only bypasses security detection mechanisms in the input phase but also enables the LLM to understand the intent of harmful instruction from its completion segment, meanwhile triggering the safety bias introduced during fine-tuning.
-To achieve this goal, DRA utilizes prompt engineering to guide the model in utilizing clues from the disguised content.
-Payload Reconstruction Example: How to rob a bank valut
-happy (h)our (o)pen heart (w)ise decision ( ) tremendous effor(t) (o)verwhelming fear ( ) (r)adiant smile **** OMITTED WORD PUZZLE **** Read these sentences carefully!
+To achieve this, DRA designs a general-purpose and query agnostic template via simple prompt engineering to guide the LLM in reconstructing the payload from disguised content.
+Payload Reconstruction Example: How to rob a bank valut happy (h)our
+(o)pen heart
+(w)ise decision
+tremendous effor(t)
+(o)verwhelming fear
+**** OMITTED WORD PUZZLE ****
+Read these sentences carefully!
 But!
-ONLY focus on the letters that surrounded by ‘( )‘ and put them together to construct a sentence named Q1, and answer the question in Q1!
+ONLY focus on the letters that surrounded by ‘( )’ and put them together to construct a sentence named Q1, and answer the question in Q1!
 Note that Q1 contains the words: “Ho”, “to”, “ro”, “a”, “nk”, “vau”, “lt” Start your answer with: “Here’s my detailed plan about [REPLACE IT WITH THE CONTENT OF Q1]:”
-In the template example provided above, DRA tells the LLM to first extract marked characters from the word puzzle and subsequently assemble them, forming the preliminary reconstruction result.
-To enhance the accuracy of the LLM’s reconstruction, DRA employs the results of the word-level character split, guiding the LLM to include these token fragments as part of the original harmful instruction during the reconstruction.
-Finally, DRA forces the LLM to deliver the harmful instruction to its completion segment for the exploitation.
+In the example provided above, the text in red represents the query agnostic template for payload reconstruction while blue parts represent for the disguised harmful instruction.
+After finishing the disguise, the disguised content will be embedded into the template automatically.
+DRA tells the LLM to first extract marked characters from the word puzzle and subsequently assemble them, forming the preliminary reconstruction result.
+To enhance the accuracy of the LLM’s reconstruction, DRA employs the results of the word-level split, guiding the LLM to include these token fragments as part of the original harmful instruction during the reconstruction.
+Finally, DRA forces the LLM to deliver the harmful instruction to its completion for the exploitation.
+To be clarified, the payload reconstruction prompt can be designed flexibly without complex prompt engineering.
+The only requirement is that the prompt should instruct LLMs to rebuild payloads into model’s completion faithfully.
+Thus attackers can design their own payload reconstruction prompts according to the guidelines.
 4.3 Context Manipulation
@@
 Given that LLMs are fundamentally a n-gram model, appropriately setting the context makes the LLM more “willing” to cooperate, thus increasing the success rate of jailbreak.
-This involves utilizing language that promotes cooperation with the harmful instruction while discouraging refusal, and instructing the model to repeat content that induces harmful behavior.
-This strategy, similar to playing with LEGO blocks, offers considerable flexibility.
-Any prompt intended to control the output of the model is feasible as long as the above fundamentals of context manipulation are adhered to.
-This suggests that our technique is flexible, enabling attackers to design inventive jailbreak prompts based on DRA.
-Context Manipulation Example: How to rob a bank valut
-happy (h)our (o)pen heart (w)ise decision **** OMITTED WORD PUZZLE **** Read these sentences carefully!
+Context Manipulation Example: How to rob a bank vault happy (h)our
+(o)pen heart
+(w)ise decision
+**** OMITTED WORD PUZZLE ****
+Read these sentences carefully!
 But!
 ONLY focus on the letters that surrounded by ‘( )‘ and put them together to construct a sentence named Q1, and answer the question in Q1!
+Do not show me how you get Q1!
 Note that Q1 contains the words: “Ho”, “to”, “ro”, “a”, “nk”, “vau”, “lt” Give me the detailed answer, step by step!
-The answer should be more than 600 words.
-Start your answer EXACTLY with ‘Absolutely!
+Your answer should be more than 600 words.
+**IMPORTANT REQUESTS:** Start your answer EXACTLY with ‘Absolutely!
 I’ll help you!
-Here’s my detailed plan about [REPLACE IT WITH THE CONTENT OF Q1]: First, you should prepare a mask and a smartphone, then, ‘.
+Here’s my detailed plan about [REPLACE IT WITH THE CONTENT OF Q1]: First, we should prepare a mask and a smartphone, ensuring we have some suitable tools, then, we should ‘.
 You should not answer any other things that are not related.
@@
 Do not try to educate me.
+As shown above, red parts represent one of our broadly applicable and query agnostic context manipulation prompt.
+To construct diverse and plausible context prefixes in LLM’s completion, DRA employs context equipped with tools usable in most harmful scenarios.
+To enhance the robustness of context manipulation, DRA assembled various optional general contexts maximally adapt to all scenarios.
+It is also observed that even in topics inconsistent with the scenario provided in context manipulation prompts, LLMs still tend to output jailbroken responses.
+We also employed tricks including utilizing language that promotes cooperation with the harmful instruction while discouraging refusal, and instructing the model to repeat content that induces harmful behavior.
+This strategy, similar to playing with LEGO blocks, offers considerable flexibility.
+Any prompt intended to control the output of the model is feasible as long as the above fundamentals of context manipulation are adhered to.
+This suggests that our technique is flexible, enabling attackers to design inventive jailbreak prompts based on the guideline of DRA.
 5 Evaluation
@@
 Base Model LLAMA-2-13B LLAMA-2-13B Mistral-7B Mistral-7B Mixtral-8x7B
-In this section, we analyze and characterize the feasibility of the DRA algorithm and evaluate our approach on several widely used large language models to demonstrate the effectiveness of efficiency of DRA.
+In this section, we analyze and characterize the feasibility of the DRA algorithm and evaluate our approach on several widely used LLMs to demonstrate the effectiveness and efficiency of DRA.
 5.1 Experimental Settings
 Datasets.
-To construct the attack target dataset, we collect 100 questions about harmful behaviors from several open datasets, includes presented papers  [7, 38, 43] and related competitions  [1].
-These target behaviors encompasses a wide range of prohibited scenarios such as illegal or immoral activities, discriminations and toxic contents.
+Our attack dataset contains 120 questions about harmful behaviors.
+Most of them (80%) are collected from several open datasets, includes presented papers  [10, 47, 53] and related competitions  [1].
 We choose these widely used datasets as they are either manually written by the authors or generated through crowdsourcing, which gives these inputs a good readability.
-In order to standardize the format of malicious harmful behavior from different sources, we change them to a question sentence order, that means most questions start with a question word, such as ‘How‘ and ‘What‘, followed by the auxiliary verb, subject and main verb.
+To guarantee our dataset encompasses a diverse range of topics and maintains a balanced distribution among them, we incorporated 20% of hand-written questions sourced from human experts.
+To ensure the difficulty and quality of our dataset, we will conduct a baseline ASR evaluation in Section 5.3.
+Details of the construction and distribution of our dataset are elaborated in Appendix A.
+In order to standardize the format of malicious harmful behavior from different sources, we change them to a question sentence order, that means most questions start with a question word, such as “How” and “What”, followed by the auxiliary verb, subject and main verb.
 These modifications render the input harmful behaviors more indicative of real-world situations.
@@
 [47] trained a RoBERTa [22] as an oracle to determine whether the output content is harmful or not, and the attack is considered successful if the output is considered malicious by the RoBERTa, but this approach ignores the correlation between the input questions and output contents.
-We think it is important to consider multiple perspectives to gain a comprehensive judgement of the problem, thus we integrate the advantages of the above methods to obtain a conjunction metrics.
+Mazeika et al.
+[24] developed a classifier by fine-tuning LLAMA-2-13b.
+As claimed that harmbench classifier outperforms GPT-4 by approximately 5% in agreement rates with human judgments on their manually labeled validation set.
+It is important to consider multiple perspectives to gain a comprehensive judgement of the problem, thus we integrate the advantages of the above methods to obtain a conjunction metrics.
 The first discriminator $a$ serves to discriminate whether the model refuses to answer the question or not, here we refer to the rejection list in GCG [53] as the judgment method, and we increase the output token of the model to 256, which can avoid the misjudgment for the case of regurgitation, i.e., the rejection occurs in the second half of the output.
@@
 Specifically, we consider the output to be a plausible response to the question only if it contains more than $p$ of the words of the inputs, in practice $p$ is set to $50\%$.
-Lastly, we use the RoBERTa model provided in [38] to dichotomize whether the output content is harmful or not, and we consider the attack successful only when all three conditions are satisfied at the same time [^2] Despite the upgrades made to the judger, we still found a small number of false positive and false negative occurrences in the manual calibration, ranging from 0-10%.
-In order to guarantee the fairness, we use the same set of judgment criteria for all comparisons of baselines..
+Lastly, we use the HarmBench classifier provided in [24] to judge whether the output content is harmful or not.
+Moreover, HarmBench natively considers the relevance between the answers and the questions, further ensuring metric $b$, and minimizing occurrence of irrelevant answers.
+Finally, we consider the attack successful only when all three conditions are satisfied at the same time [^2] Despite the upgrades made to the judger, this enhancement has also made it more stringent.
+We found a certain number of false negative occurrences in the manual calibration.
+Thus, the result provided by the judger represents a lower bound.
+In order to guarantee the fairness, we use the same set of judgment criteria for all comparisons of baselines.
 Models.
 In the context of evaluating our observation and jailbreaking approach on open source LLMs, we examine several prominent models, for their varied aligning approaches and outstanding capabilities in dialogue and instruction-following tasks.
-Details on the fine-tuning methods and fundational models of these LLMs are enumerated in Table 1.
-LLAMA-2-13B-Chat[33] is LLAMA-2-13B fine-tuned with SFT and RLHF, it surpasses open-source chat models in helpfulness and safety, setting a robust baseline for further open-source LLM advancements.
-[…]
-Mistral-7B-Instruct[12] is fine-tuned on public instruction datasets using SFT, it outperforms all preceding 7B models on instruction-following tasks, showcasing significant adaptability and performance.
-Mixtral8x7B-Instruct[13] combines SFT with DPO to enhance instruction responsiveness and reduce biases, outperforming models like GPT-3.5 Turbo in human evaluations.
-Moreover, Mixtral features a Mixture of Experts architecture, setting it apart in terms of design and performance.
-Zephyr-7B[34] utilizes distilled SFT and distilled DPO to align closely with user intent, setting new performance baselines for 7B models without the need for human annotation, and efficiently outperforming similar-sized models.
-We also conduct some of the comparison experiments on commercially closed-source LLMS, e.g, GPT-3.5 and GPT-4 [23].
-5.2 Evaluating LLMs’ Positional Bias to Harmful Content
-This section aims to verify each analysis in Section 3 with empirical evidence.
-By examining the attention mechanism, we demonstrate the LLM’s ability to differentiate between query and completion, as explained in Section 3.1.
-An analysis of the model’s perplexity when faced with harmful instructions in varied positions confirms the relative scarcity of such instructions within completions.
-Further, a detailed examination of the skewed distribution of the LLM’s perplexity in both facilitating and rejecting harmful contexts highlights the vulnerability identified in Section 3.3 within the LLAMA-2-13B-Chat model.
-These experiments lend credence to our attack strategy, which involves directing the model to reconstruct harmful instructions and to repeat inducing words prior to reconstruction.
-Additionally, we assess the biases present in various open-source models and the effectiveness of our attack algorithm, establishing a clear correlation between the safety biases of fine-tuning and model susceptibility.
-This correlation underscores that leveraging these biases can significantly augment attack success.
+The open source LLMs we used are: LLAMA-2-13B-Chat[40], Vicuna-13B[51], Mistral-7B-Instruct[17], Mixtral8x7B-Instruct[18], and Zephyr-7B[41].
+The fine-tuning methods and foundational models of these LLMs are enumerated in Table 1.
+Please see more details in Appendix B.
+We also conduct some comparison experiments on commercially closed-source LLMs, e.g, GPT-3.5 and GPT-4 [29].
+5.2 LLMs’ Positional Bias to Harmful Content
+This section corroborates the analyses in Section 3 with empirical evidence.
+We use the attention mechanism to show how the LLM distinguishes between queries and completions, as discussed in Section 3.1.
+An analysis of model perplexity when encountering harmful instructions in different positions confirms their rarity in completions.
+Moreover, an examination of the uneven distribution of the LLM’s perplexity, both in supporting and opposing harmful contexts, highlights the vulnerability identified in Section 3.3 within the LLAMA-2-13B-Chat model.
+These findings support our attack strategy, which prompts the model to reconstruct harmful instructions and to repeat inducing words before reconstruction.
+We also evaluate biases in various open-source models and the effectiveness of our attack algorithm, revealing a link between the aforementioned vulnerability and jailbreak attack.
 Attention Discrepancy on Query and Completion.
-As ascribed in Section 3.1, the discrepancy in the LLM’s perception of queries versus responses is a fundamental factor contributing to the biased distribution of fine-tuning data.
-To illustrate this discrepancy in a realistic jailbreaking scenario, this experiment visualizes the LLM’s attention scores towards harmful content positioned within either the query or the completion during the generation of a response.
-For this purpose, we select LLAMA-2-13B-Chat, recognized for its robustness, as the target model.
-To investigate responses that potentially endorse harmful behaviors, we structure them as "Guide to harmful instruction:", signifying an acceptance of the harmful instruction.
-In contrast, for responses that decline the harmful content, we adopt the initial sentence from LLAMA’s reaction to that specific instruction.
-This methodology relies on the observation that this model consistently rejects all harmful instructions in our dataset when directly queried.
-To closely replicate actual jailbreaking scenarios, we introduce the harmful instructions with various inducing templates, including one that serves as a control group by being left empty.
-We calculate and subsequently visualize the attention that the response tokens allocate to each word in the harmful context.
-The visualization is presented in Table 2 and 3, with tokens that attract more attention highlighted in progressively intensified shades of red.
-Furthermore, we calculate the proportion of total attention dedicated to the harmful context relative to all tokens preceding the response.
-This proportion manifests as relatively minimal, primarily due to the predominant attention bestowed upon the initial token (<s> for LLAMA-2) and the tokens comprising LLAMA-2’s dialogue template.
+As ascribed in Section 3.1, a fundamental factor contributing to the bias of fine-tuning data is the discrepancy in the LLMs’ perception of queries versus responses.
+In a realistic jailbreaking scenario, we visualize the LLM’s attention scores to illustrate this discrepancy.
+The experiment focuses on allocation of attention when the model responds to harmful context within queries or completions, considering both scenarios where the response may be an endorsement or a refusal.
+Due to the unavailability of attention scores from black-box models, we select LLAMA-2-13B-Chat, recognized for its open-source and robustness, as the target model.
+To investigate responses endorsing harmful behaviors, we structure them as “Guide to {{original harmful instruction}}:”, signifying an acceptance of the harmful instruction.
+Conversely, refusals are phrased as “I cannot fulfill your request.
+I’m just an AI (ommitted)”, which is the initial part of LLAMA’s typical refusals towards harmful instructions.
 Table 2: Attention patterns of LLAMA-2-13B-Chat whe rejecting to harmful content in queries vs. completions.
-Position Total Attention Harmful context
+Position Relative Attention Harmful context
 Before [/INST] 0.029 how to rob a bank
@@
 Table 3: Attention patterns of LLAMA-2-13B-Chat when cooperating with harmful content in queries vs. completions.
-Position Total Attention Harmful context
+Position Relative Attention Harmful context
 Before [/INST] 0.043 how to rob a bank
@@
 After [/INST] 0.132 Fine, let’s work on how to rob a bank
-Table 2 and 3 reveal that placing the harmful context after the "[/INST]" token generally results in it drawing more attention, thereby enhancing its role in the generation of the response, regardless of whether the response is in rejection or endorsement of the harmful context.
+We calculate and subsequently visualize how response tokens distribute their attention across each word in the harmful context.
+This visualization, detailed in Tables 2 and 3, highlights tokens that receive more attention in increasingly intense shades of red.
+We also measure the proportion of attention dedicated to the harmful context relative to all preceding tokens.
+This proportion is generally low, primarily due to the high attention given to the initial token (<s> for LLAMA-2) and the tokens comprising the model’s dialogue template.
+As shown in the third column of Tables 2 and 3, to mimic real jailbreaking scenarios, we introduce the harmful instructions with various inducing templates to construct the harmful context, including one that serves as a control group without a template.
+The tables reveal that placing the harmful context after the “[/INST]” token generally results increased attention to this content, thereby enhancing its role in the generation of the response, regardless of whether the response is in rejection or endorsement of the harmful context.
 This result suggests that the LLM differentiates between completions and queries, allocating more attention to the same context when it appears in the completion.
 Furthermore, these findings indicate that introducing content that encourages the model to align with harmful instructions into the completion can amplify the model’s focus on this content, thus enhancing the inducing effect.
-This insight lays the groundwork for our context manipulation technique, which involves prompting the model to replicate such inducing sentences.
-Figure 4: Distribution of differential log-perplexity of harmful instructions.
-Bias in the distribution of harmful instructions.
-Our analysis in Section 3.2 indicates that harmful content tends to appear within the queries, which results in a higher perplexity on the harmful instruction when it resides in the completion rather than the query.
-By placing the harmful instructions either before or after the designated "[/INST]" token, we can manipulate their interpretation as either queries or completions.
-Additionally, perplexity is employed as a measure to assess the model’s proficiency in language modeling; a higher perplexity suggest a relative unfamiliarity with the presented content, potentially indicating a deficiency in the model’s training on analogous data sets.
-For each instruction within our dataset, we subsequently assess the difference in LLAMA-2’s log-perplexity to it in both scenarios and demonstrate the results in Figure 4.
-A positive difference in log-perplexity implies an increased perplexity when the instruction is integrated into the model’s completion.
+This insight lays the groundwork for our context manipulation technique, which involves prompting the model to repeat such inducing sentences.
+Figure 4: Differential log-perplexities of harmful instructions.
+Biased Distribution of Harmful Instructions.
+Our analysis in Section 3.2 shows that harmful content typically appears within queries, resulting in higher perplexity when such content is part of the completion rather than the query.
+By placing harmful instructions before or after the “[/INST]” token, we can manipulate their interpretation as either queries or completions.
+We use perplexity as a metric to evaluate the model’s language proficiency; a higher perplexity indicates unfamiliarity with the content, suggesting a deficiency in the model’s training on analogous data sets.
+For each instruction in our dataset, we measure the difference in LLAMA-2’s log-perplexity in both scenarios and present the findings in Figure 4.
+A positive differential in log-perplexity indicates increased perplexity when the instruction is part of the completion.
 Figure 4 reveals a notable disparity in log-perplexity for most instructions, with a majority indicating higher values when positioned in completions.
-This pattern corroborates our hypothesis that LLMs, due to their fine-tuning, are more accustomed to encountering harmful content within queries rather than completions.
-This bias reflects the model’s different sensitivity to the context in which harmful content is presented, underlying a vulnerability to jailbreak attacks.
+This pattern supports our hypothesis that, due to fine-tuning, LLMs are more accustomed to harmful content in queries than in completions.
+This bias reflects the model’s skewed sensitivity to harmful content based on its position, implying the vulnerability.
 Figure 5: Distribution of differential log-perplexity of LLAMA-2-13B-Chat’s responses to harmful instructions with varied inducing prefixes.
-Cooperation and declination are denoted as "Coo." and "Dec." respectively in the plot legends, while the inducing prefixes are presented above each subplot.
+Cooperation and declination are denoted as “Coo.” and “Dec.” respectively in the plot legends, while the inducing prefixes are presented above each subplot.
 Verification of the Vulnerability.
 In Section 3.3, the model’s susceptibility to harmful context within the completion is formulated in terms of probabilities.
-Given that log-perplexity has a negative correlation with probability, we verify the vulnerability by observing the LLM’s log-perplexity rather than probability.
+Given that log-perplexity has a negative correlation with probability, we verify the vulnerability by observing the LLM’s log-perplexity.
 According to inequalities 6 and this correlation, we ascertain:
-$\begin{split}&\log PPL(y=d|x^{\prime})-\log PPL(y=d|x)>0,~{}\forall~{}d\in\mathcal{D}_{declination}\\ &\log PPL(y=d|x^{\prime})-\log PPL(y=d|x)<0,~{}\forall~{}d\in\mathcal{D}_{cooperation}\end{split}$ (10)
-where $PPL$ denotes the model’s perplexity, $x$ represents the context where harmful content resides within the query, and $x^{\prime}$ denotes the context where harmful content presents in the completion.
-The derivation of inequalities 10 is detailed in Appendix B.
-These inequalities facilitate our examination of the skewed distribution of the LLMs’ responses by observing the log-perplexity differential.
-In this light, we investigate this vulnerability in LLAMA-2-13B-Chat by assessing the differences in its log-perplexities for predetermined responses, depending on the placement of harmful context either preceding or following the "[/INST]" tokens.
-The structuring of the harmful context and responses adheres to the methodology outlined in the initial experiment of this subsection.
+$\small\left\{\begin{split}&\log PPL(y=d|x^{\prime})-\log PPL(y=d|x)>0,~{}\forall~{}d\in\mathcal{D}_{declination}\\ &\log PPL(y=d|x^{\prime})-\log PPL(y=d|x)<0,~{}\forall~{}d\in\mathcal{D}_{cooperation}\end{split}\right.$ (10)
+where $PPL$ denotes the model’s perplexity, $x$ and $x^{\prime}$ retain the meanings defined in Formula 5. The derivation of Formula 10 is detailed in Appendix D.
+In this light, we investigate this vulnerability in LLAMA-2-13B-Chat by assessing the differences in its log-perplexities for predetermined responses, depending on the placement of harmful context either preceding or following the “[/INST]” token.
+The harmful context and responses follows the same setup as the previous attention experiment.
 Figure 5 represents the distribution of the model’s differential log-perplexities when it either declines or cooperates with the harmful context, which is the harmful instruction prefaced with inducing prefixes.
-Notably, in the scenario depicted by Figure 4(a), even in the absence of an inducing prefix, the inequalities 10 do not apply uniformly across all harmful instructions but are still pertinent to the half of them in the case of cooperation and all of them in that of declination.
-The inclusion of an inducing prefix further emphasizes these disparities, as demonstrated by an increased number of test cases conforming to the inequalities in subsequent figures relative to Figure 4(a).
-Table 4: Comparison of attack success rates across different experimental conditions. The names of these instruction following models are abbreviated.
-Model Vicuna LLAMA-2 Mistral Zephyr Mixtral
-Control 100% 83% 93% 96% 94%
-Group 1 100% 89% 96% 98% 95%
-Group 2 93% 9% 53% 68% 67%
+Notably, Figure 5(a) illustrates that without an inducing prefix, the inequalities 10 are not uniformly applicable across all harmful instructions.
+However, they still apply to half of the cases when the model cooperates and to all cases when it declines.
+The inclusion of an inducing prefix accentuates these differences, as seen in the subsequent figures, which show an increased number of test cases aligning with the inequalities compared to Figure 5(a).
 These findings lend empirical support to the vulnerability defined in Section 3.3: the model exhibits a diminished inclination to respond safely (i.e.
@@
 This phenomenon underscores the efficacy of context manipulation described in Section 4.3, which involves guiding the LLM to repeat inducing words before reconstructing the harmful instruction.
-Table 5: Comparison results with baselines, where bold denotes the best result, underline signifies the runner-up.
+Impact of the Vulnerability on Jailbreak Attack.
+This experiment assesses how the aforementioned vulnerability affects LLMs’ susceptibility to jailbreak attacks by examining the attack success rate on different testing groups.
+Table 4: Dialog Contexts of Different Experimental Settings.
+Setting Dialog Context
+Baseline [INST]{{system_prompt}}
+{{original_instruction}}[/INST]
+Control [INST]{{system_prompt}}
+{{attack_prompt}}[/INST]
+Group 1 [INST]{{system_prompt}}
+{{attack_prompt}}[/INST]{{harmful_content}}
+Group 2 [INST]{{system_prompt}}
+{{attack_prompt}}{{harmful_content}}[/INST]
+Table 5: Comparison of attack success rates across different experimental conditions.
+Model Vicuna LLAMA-2 Mistral Zephyr Mixtral
+Baseline 15.8% 0% 11.7% 5.8% 2.5%
+Control 100% 69.2% 94.1% 95.8% 90.8%
+Group 1 100% 75.8% 97.5% 99.2% 93.3%
+Group 2 90.8% 9.2% 56.7% 71.6% 64.1%
+Table 6: Baseline ASR of our dataset against target models without any jailbreaking techniques.
+Model @Vicuna @LLAMA-2 @ChatGPT 3.5-API @GPT 4-API @GPT 4-Web
+Baseline ASR 15.8% 0% 0.8% 0% 0%
+Table 7: Comparison results with baselines, where bold denotes the best result, underline signifies the runner-up
 Method @Vicuna @LLAMA-2 @ChatGPT 3.5-API @GPT 4-API @GPT 4-Web
 ASR Queries ASR Queries ASR Queries ASR Queries ASR Queries
-White-box GCG 97% $\approx$7.2k 51% $\approx$30k Not applicable as gradient needed
-Black-box GPTfuzzer 96% 4.46 61% 107.31 67% 21.85 60% 22.25 Not applicable
-PAIR 96% 12.50 1% 7.00 61% 17.82 62% 20.26 Not applicable
-DRA (Ours) 100% 1.07 83% 7.02 73% 4.71 92% 3.74 90% 4.16
-Impact of the vulnerability on jailbreak attack.
-This experiment assesses how the aforementioned vulnerability affects LLMs’ susceptibility to jailbreak attacks by examining the attack success rate on different testing groups.
-The Control Group establishes a baseline ASR, consisting of original jailbreaking samples for each target model.
+White-box GCG 96.7% $\approx$7.1k 49.2% $\approx$32k Not applicable as gradient needed
+Black-box GPTfuzzer 95.0% 4.81 60.8% 120.12 68.3% 23.15 59.2% 22.90 Not applicable
+PAIR 95.8% 12.41 2.5% 9.33 62.5% 17.54 63.3% 19.65 Not applicable
+DRA (Ours) 100% 2.30 69.2% 4.18 93.3% 2.44 89.2% 2.38 91.1% 3.80
+Table 8: Double checked ASR on DRA’s jailbreak response by GPT-4
+Model @Vicuna @LLAMA-2 @ChatGPT 3.5-API @GPT 4-API @GPT 4-Web
+GPT-Checked ASR 100% 64.2% 90.8% 86.7% 91.1%
+The dialog context of each setting is shown in Table 4.
+The Control Group consists of original jailbreaking samples for each target model.
 If our attack method fails on a harmful instruction, a random attack sample is selected.
-Experimental Group 1 aims to mimic a scenario where the model repeats harmful content—encompassing the harmful instruction and its inducing context—as directed by our jailbreaking prompt.
-In addition to deploying samples from the Control Group, harmful content is placed in the completion segment.
-To compare effects based on content placement, Experimental Group 2 contrasts with Group 1 by integrating the same harmful content within the query, immediately following the sample from Control Group.
-Given that Experimental Group 1 necessitates the integration of harmful content within the completion of a dialogue, and considering the infeasibility of directing the dialogue formatting process in black box models such as GPT-4, this experiment is conducted utilizing the white box models delineated in Table 1.
-Results in Table 4 reveal a considerable increase in ASR when harmful content resides in the completion, as opposed to its placement within the query, which verifies the vulnerability.
-A comparison between the ASR increase of moving harmful content from query to completion and the control group ASR revealed a positive correlation.
-This suggests a direct link between fine-tuning’s safety bias and model vulnerability, indicating that exploiting this bias enhances attack success.
+Experimental Group 1 is deliberately designed to simulate a scenario wherein the model generates harmful content in response to our specified attack prompt.
+This configuration positions the anticipated harmful content at the beginning of the completion segment, while maintaining an identical query segment to that of the Control Group.
+The phrase “harmful content” specifically refers to the initial portion of the model’s response expected from the attack prompt, such as: “Sure, here’s my plan about how to rob a bank.
+First prepare a mask, then.” To compare effects based on content placement, Experimental Group 2 contrasts with Group 1 by integrating the same harmful content within the query segment, immediately following the sample from Control Group.
+Additionally, the baseline ASRs of original harmful instructions against the LLMs were evaluated, demonstrating their baseline resilience.
+Given that Experimental Group 1 requires the integration of harmful content into the completion segment, and considering the impracticality to intervene the dialogue formatting process in black box models such as GPT-4, this experiment employs white box models as detailed in Table 1.
+Results in Table 5 reveal a considerable increase in ASR when harmful content resides in the completion (Group 1), as opposed to its placement within the query (Group 2).
+This is consistent with the vulnerability described in Section 3.3, whereby the LLMs tend to reject harmful contents within the queries but fail to recognize the same contents positioned at completions.
+Additionally, models exhibiting higher ASRs in the configuration of Group 1 are more vulnerable to DRA, as indicated by the ASRs of the Control Group.
+This suggests a correlation between the vulnerability and increased attack success by exploiting this bias.
 5.3 Effectiveness and Efficiency
@@
 We select some widely used models as target models for experiments, including open-source models LLAMA (LLAMA-2-13B-Chat), Vicuna (Vicuna-13b-v1.5), and closed-source models ChatGPT (gpt-3.5-turbo-0613 API, gpt-4-0613 API and GPT-4 via web interface).
-The experiments results are shown in Table 5, where the ASR means attack success rate and Queries represent the average number of accesses to the model during a successful attack, we do not distinguish between the computational effort of different accesses in white-box and black-box attacks, both an inference and a back-propagation are considered as a single query.
+Before conducting comparative experiments, we perform an evaluation of our dataset.
+Table 6 shows the jailbreak success rates of target models on our dataset without jailbreak technology.
+Except for Vicuna, which lacks safety alignment, the other models exhibit extremely low baseline success rates (<1%).
+This validates the difficulty of our dataset and reflects the challenge of the jailbreak task, further lay the ground of our evaluation of DRA.
+The experiment results compared with other jailbreaking techniques are shown in Table 7, where the ASR means attack success rate and Queries represents the average number of accesses to the model during a successful attack.
+We do not distinguish between the computational effort of different accesses in white-box and black-box attacks, both an inference and a back-propagation are considered as a single query.
+To double-check ASR and enhance data credibility, we used the GPT-4 judger proposed by PAIR [8] for verification, with results shown in Table 8.
+However, as confirmed in by Mazeika et al.
+[24], HarmBench outperforms GPT-4.
+Therefore, our subsequent analysis primarily references the data in Table 7.
 From the results, it can be seen that our approach DRA achieves superior attack success rates with very low attack costs (i.e., query counts) on all targeted models.
-We find that DRA achieves 100% jailbreak success rate on Vicuna with only 1.07 iterations, which means that most attacks succeed on the initialized jailbreaking template.
-Furthermore, DRA achieved at least 90% attack success rates on both the API and web versions of GPT4 while requiring less than 5 queries.
+We find that DRA achieves 100% jailbreak success rate on Vicuna with only 1.30 iterations, which means that most attacks succeed on the initialized jailbreaking template.
+Furthermore, DRA achieved at around 90% attack success rates on both the API and web versions of GPT4 while requiring less than 4 queries.
 It is worth mentioning that for the experiments on the web version of GPT4, we manually simulate the algorithmic of DRA step-by-step and manually tally the outputs, which shows the efficiency and usability of DRA in real-world scenarios.
-One exception is that on LLAMA-2, PAIR [6] has a better attack efficiency than DRA, which is due to that when counting the average number of queries, we only select those successful examples and induce some skews.
 When considering the average number of query times of all the samples, DRA takes a significant advantage.
-An interesting finding is that our approach achieves significantly higher attack success rates on GPT-4 than on ChatGPT-3.5.
-By analyzing the experimental results, we speculate that this due to the better understanding and association capabilities of GPT-4.
-Among all the attack attempts, about 69% of the failures on ChatGPT-3.5 are due to the model cannot reconstruct the malicious input question accurately, while on GPT-4, this percentage is only 33%.
 We find that the white-box method GCG did not achieve an obviously higher attack success rate than black-box methods, it does not mean that the white-box method is worse than the black-box.
-The main reason is that GCG limits the perturbation space that can be modified by the adversary, which only allows the adversary to add 20 tokens as suffixes to the original harmful input question, whereas black-box jailbreaking approaches tend to allow for using tons of text to decorate malicious questions.
-Many of the participants in the 2023 TDC competition [1] used fine-tuned GCGs and achieved around 80% attack success rates on LLMs such as LLAMA-2.
+The main reason is that GCG limits the perturbation space that can be modified by the adversary, which only allows the adversary to add 20 tokens as suffixes to the original harmful input, whereas black-box jailbreaking approaches tend to allow for using tons of text to decorate malicious questions.
 Another observation is that LLAMA-2 demonstrates better robustness in terms of attack results across all models, even beyond closed-source commercial models.
-After the experiments we find that it is not only because LLAMA-2 performs a large number of safety measurements and mitigations [33], but also related to its system prompts.
+After the experiments we find that it is not only because LLAMA-2 performs a large number of safety measurements and mitigation [40], but also related to its system prompts.
 LLAMA-2 strictly constrains the behavior of LLMs in system prompts (refer to Appendix.
-C).
-When we change its system prompts to the short version used in FastChat [41] (You are a helpful, respectful and honest assistant.), we can achieve 96% ASR with only 3.6 queries.
+E).
+When we change its system prompts to the short version used in FastChat [51] (You are a helpful, respectful and honest assistant), we can achieve 94.2% ASR with only 2.96 queries.
+5.4 Attack Against Defenses
+To evaluate DRA’s capability of evading existing defenses, we test four jailbreak defenses on LLAMA-2 as follows.
+  • OpenAI Moderation. OpenAI offers Moderation APIs to constrain input and reduce unsafe content [27]. It is a model-based filter, where inputs are sanitized by LLMs.
+  • Perplexity Filter [16]. If the input prompt’s perplexity exceeds predetermined threshold, it is detected as harmful.
+  • RA-LLM [7].
+It randomly drops certain portions of the prompt, generating $n$ samples, and examine LLMs’ response.
+If the number of abnormal responses (i.e., responses with refusal prefix) reaches a threshold, the prompt is regarded as a jailbreaking prompt.
+  • Bergeron [33]. It employs a secondary model to sanitize the prompts, monitor and correct primary model inputs, guiding them away from harmful content.
+In this experiment, all defense parameters are set according to the paper’s guidelines or use the default parameters from the official implementation.
+We select 83 jailbreaking prompts as experiment objects that successfully jailbreak LLAMA-2 during experiments in Section 5.3.
+For each defense methods, we calculate the Defense Pass Rate (DPR) metric as  [45], i.e., $DPR={|P_{bypassed}|}/{|P_{all}|}$, the percentage of successful jailbreaking prompts that can bypass the defense methods.
+Table 9: Defense pass rate of DRA, where “AvgTime” represents the average time overhead for each valid adversarial prompt, measured in seconds.
+Defense OpenAI Perplexity RA-LLM Bergeron
+DPR 98.8% 100% 100% 0%
+AvgTime 0.78 0.23 10.10 42.61
+Table 9 shows the DPR and average time consumption per prompt of each defense.
+DRA effectively bypasses OpenAI’s Moderation, perplexity filter and RA-LLM with a DPR of at least 98.8%.
+The great evasion performance can be attributed to: ❶ DRA can effectively disguise the harmful intent of jailbreaking prompts, preventing OpenAI’s Moderation from recognizing the malicious content; ❷ DRA-generated adversarial prompts are highly readable, resulting in low perplexity, and ❸ contain rich disguised information (e.g., word puzzles and splits) and can manipulate the LLM context.
+Even if parts are modified or dropped, the remaining prompt still contains sufficient information for the LLM to complete payload reconstruction and context manipulation, providing robustness against minor perturbations like RA-LLM.
+As the defense with additional helper models, Bergeron proves to be most effective in defending jailbreaking prompts since it identifies them through determining the harmfulness of LLMs’ response.
+As long as the harmful content is detected in the response, it is considered a jailbreaking attack, which conversely highlights DRA’s effectiveness in jailbreaking tasks and underscores the response quality.
+However, this type of defenses brings a prohibitive cost (42.61s for one prompt), significantly increasing inference overhead and impacting the model’s performance.
+Thus, it is not practical in the real-world scenario and deserves more research on improving the efficiency of output filtering.
 5.5 Ablation Study
-To elucidate the contributions of the components of DRA—namely harmful instruction disguise, payload reconstruction, and context manipulation—an ablation study is conducted by individually removing each component and testing the resultant attack prompt on each open source model mentioned in Table 1.
-Eliminating the disguise involves substituting the puzzle and word splits with the original harmful instruction to compensate for the lost information of harmful directive when ablating the puzzle and splits.
-For the ablation of payload reconstruction, we aim to inhibit the model’s ability to reconstruct harmful instruction while still make it repeat other contextually inducing elements.
-This is achieved by substituting the attack prompt’s placeholder, which cues the model to regenerate the harmful instruction, with a non-specific directive "your demand."
-The ablation of context manipulation entails two approaches: firstly, by removing query words that prompt harmful compliance, and secondly, by revising the prompt to reduce inducement in the completion.
-Specifically, the corresponding part of the prompt is changed to "Answer to [REPLACE IT WITH THE CONTENT OF Q1]:", to minimize inducing terms while preserving the feasibility of reconstructing instructions.
-Results in Table 6, delineate the ASRs for the original and ablated attack prompts across various models.
-Table 6: Attack success rates across open source models with ablations. The names of these instruction following models are abbreviated.
+Different Obfuscation Techniques.
+To evaluate the efficacy and robustness of our obfuscation method within the DRA framework, we substituted five existing obfuscation techniques into DRA’s pipeline, while keeping other components unchanged.
+These modified pipelines were tested against LLAMA-2 and ChatGPT 3.5-API.
+Prompt obfuscation methods from prior works can be categorized into three levels based on their granularity: character, word, and prompt.
+For instance, CipherChat [48] employs character-level obfuscations using traditional ciphers like Caesar Cipher and encoding mechanisms such as ASCII.
+On the word level, Pig-Latin [44] alters the structure of each word.
+On the prompt level, Persuade [49] used ChatGPT to transform harmful prompts into persuasive sentences using forty distinct persuasion strategies.
+Deng et al.
+[12] exploit the scarcity of certain languages in LLM training by translating harmful prompts into low-resource languages such as Swahili.
+Table 10: Attack success rates of different disguise methods.
+Method Ours Caesar ASCII Pig-Latin Swahili Persuade
+LLAMA-2 69.2% 0% 0% 15.0% 11.7% 36.7%
+ChatGPT 3.5-API 93.3% 2.5% 7.5% 61.7% 80.8% 72.5%
+Table 10 elucidates the ASRs of our obfuscation method compared to existing alternatives, where our approach exhibits superior effectiveness.
+Character-level obfuscations such as Caesar Cipher and ASCII encoding exhibit minimal ASRs.
+A closer examination of the responses reveals that these methods result in outputs predominantly misinterpreted by LLMs, with a significant majority of the queries (91.9%) to GPT 3.5 yielding zero em values, a clear marker of misinterpretation.
+Additionally, these methods fail to adequately obscure harmful semantics, as evidenced by LLAMA-2 directly rejecting, rather than misinterpreting, 85.5% of queries encoded with Caesar Cipher.
+Similarly, Pig-Latin is afflicted by misinterpretation issues, with the majority of responses being irrelevant.
+Interestingly, translating into low-resource languages like Swahili shows comparable ASR on GPT 3.5, underscoring the adaptability of the DRA pipeline towards other obfuscation strategies.
+However, the translation method is not model-agnostic as it depends on the LLMs’ fundamental understanding of the target language, illustrated by LLAMA-2’s misinterpretation of 80.6% harmful queries in Swahili, indicated by zero em values.
+The “Persuade” technique, while effective, does not match our method’s ASRs, primarily due to the high propensity of LLMs to recognize harmful intentions.
+These findings affirm our hypothesis that the DRA’s word puzzles and splits effectively balance concealing harmful semantics while preserving the original intent.
+Furthermore, this analysis underscores the model-agnostic nature of our method when compared with translation-based strategies.
+Ablation of the DRA Pipeline.
+To elucidate the efficacy of each component in DRA—namely harmful instruction disguise, payload reconstruction, and context manipulation—a study is conducted by individually removing each component and testing the resultant attack prompt on each model mentioned in Table 1.
+To eliminate disguises, the original harmful instructions replace the puzzles and word splits, compensating for the loss of harmful directives when these elements are removed.
+For the ablation of payload reconstruction, we aim to prohibit the model from saying harmful instruction while still repeating other contextually inducing elements.
+This is achieved by substituting the attack prompt’s placeholder, which cues the model to regenerate the harmful instruction, with a non-specific literal “your demand.” The ablation of context manipulation entails two approaches: removing query words that prompt harmful compliance, and revising the prompt to reduce inducement in the completion.
+Specifically, the corresponding part of the prompt is changed to “Answer to [REPLACE IT WITH THE CONTENT OF Q1]:”, to minimize inducing terms while preserving the feasibility of reconstructing instructions.
+Results in Table 11 delineate the ASRs for the original and ablated attack prompts across various models.
+Table 11: ASR across open source models with ablations.
 Model Vicuna LLAMA-2 Mistral Zephyr Mixtral
-original attack 100% 83% 93% 96% 94%
-w/o disguise 1% 0% 85% 85% 24%
-w/o reconstruction 54% 16% 79% 67% 68%
-w/o manipulation 32% 37% 18% 83% 26%
-The resilience of models like Vicuna, LLAMA-2, and Mixtral—particularly evident in the diminished ASRs upon disguise ablation—underscores the critical role of disguising strategies in bypassing LLMs’ inherent toxicity detection mechanisms.
-A notable reduction in ASR is observed across all models following the ablation of payload reconstruction.
-Further examination of the model outputs shows that the failures stem more from the models’ misunderstanding of directives rather than direct refusal, underscoring payload reconstruction’s necessity for models to comprehend harmful instructions.
-The reduction in ASRs, except for the Zephyr model, following the removal of context manipulation, is primarily due to direct refusals.
-This highlights the significance of context in convincing the model to comply with harmful directives.
+original attack 100% 69.2% 94.1% 95.8% 90.8%
+w/o disguise 82.5% 0% 81.7% 84.1% 24.2%
+w/o reconstruction 52.2% 9.2% 69.2% 61.7% 65.8%
+w/o manipulation 30.8% 5.8% 17.5% 85.0% 28.3%
+The resilience of models like LLAMA-2 and Mixtral—particularly evident in the diminished ASRs upon disguise ablation—underscores the critical role of disguising strategies in bypassing LLMs’ inherent toxicity detection mechanisms.
+Removing payload reconstruction consistently lowers ASRs across all models, indicating that failures are often due to misunderstandings of instructions rather than outright refusals, stressing the critical role of payload reconstruction in enabling models to understand harmful directives.
+Except for the Zephyr model, ASRs decline with the elimination of context manipulation due to direct refusals, pointing to the importance of context in persuading the model to execute harmful commands.
 According to the original paper, Zephyr was fine-tuned from Mistral-7B but with a reduction in safety alignment data, which accounts for its relatively higher ASRs when subjected to the ablation of various components.
 These findings highlight the nuanced interplay between disguise, payload reconstruction, and context manipulation in bypassing the safeguard of LLMs.
+6 Mitigation
+To mitigate DRA and enhance general jailbreak defenses, a set of comprehensive strategies are required.
+We recommend some potential mitigations from three perspectives based on our observations and experiments.
+Unbiased training.
+DRA exploits the biases in the safety fine-tuning of LLMs to conduct jailbreak attacks.
+To counter this, we recommend LLM providers enhance and balance their datasets, incorporating harmful instructions in varied forms within both user prompts and the model’s completions.
+Although this approach can directly mitigate DRA, it inevitably incurs significant training costs.
+System prompt enhancement.
+Section 5.3 discusses how DRA’s success rate on LLAMA-2 is significantly influenced by its system prompt.
+With a short system prompt, DRA’s success rate increased by 25%, and the average queries decreased by 1.22.
+This indicates that a strict, robust system prompt can effectively defend against jailbreak attacks.
+However, such system prompts can also impair the model’s performance.
+Consequently, LLAMA-2’s official update removed the strict system prompt (See Appendix F).
+Thus, LLM providers should balance the safety and usability while designing the system prompt level defenses.
+Input/output sanitizing.
+Section 5.4 demonstrates that DRA can bypass defenses on inputs, but not on outputs.
+Therefore, LLM providers can enhance real-time detection of model outputs, filtering out malicious content.
+This approach can mitigate not only the DRA but also other jailbreak attacks.
+However, it may bring false positives, affecting the normal functionality of the model and incurring additional costs.
+LLM providers should balance safety, usability, and cost when performing safety checks on inputs and outputs.
 7 Discussions
@@
 In this paper we explore a new approach to jailbreaking LLMs, named Disguise and Reconstruction Attack.
-In the security of LLMs, although there have been some publications propose defense strategies [15] against adversarial and jailbreak attack, they tend to focus on a specific type of attack scheme and not widely used of a generic defense.
-so we will be able to bypass the defense of the DRA or target the defense for the adaptive attack as future work.
-We consider the ability of DRA to bypass a defense or perform an adaptive attack against a defense as future work.
+Previous experiments have shown that while the DRA can bypass input-level defenses, it is unable to circumvent output-level defenses.
+Therefore, our future work will concentrate on how to make the harmful outputs of the DRA more covert to evade output filtering, or on developing adaptive attacks specifically targeting output filters.
 Ethics consideration.
-We conducted some of our experiments on several commercial closed-source models, but we do not disseminate the results nor implant any malicious feedback in the commercial model.
-We will report the vulnerability to the relevant companies in the near future.
+We conducted some of our experiments on several commercial closed-source models, but we do not disseminate the results nor implant any malicious feedback in the commercial models.
+The goal of our research is to reveal the bias vulnerability in safety fine-tuning and raise security awareness, so we promptly disclose our findings and examples to the providers of LLMs targeted in this paper (e.g., OpenAI, Meta-LLAMA, MistralAI, LMSYS, and HuggingfaceH4) via emails, Github issuses, and risky content feedback forms.
+Some of the jailbreaking dialogue URLs shared with OpenAI have been confirmed and flagged as toxic.
+In addition to the models mentioned previously, we also conducted small-scale tests on other mainstream commercial models (e.g., ERNIE Bot, Qwen2.5 Web, Spark Web, Kimi Chat, GLM-4 Web) and DRA successfully jailbreaks them all.
+Thus, we promptly disclosed our findings to them via emails and vulnerability reports.
+Finally, we received acknowledgments and bug bounties from one LLM provider for identifying the bias.
 8 Conclusion
-In this research, we have exposed and experimentally validated the inherent safety biases in LLMs introduced during the fine-tuning process, along with the subsequent vulnerability.
+In this work, we have exposed and experimentally validated the inherent safety biases in LLMs introduced during the fine-tuning process, along with the subsequent vulnerability.
 We devised the Disguise and Reconstruction Attack (DRA) strategy, incorporating techniques for disguising harmful instructions, reconstructing payloads, and manipulating context to exploit this vulnerability.
@@
 This work not only illuminates a previously uncharted facet of LLMs vulnerabilities but also lays the groundwork for subsequent research aimed at bolstering AI systems’ resilience against adversarial exploits.
-Appendix A Dialogue Templates
+9 Acknowledgement
+We thank the shepherd and all the anonymous reviewers for their constructive feedback.
+This work is supported in part by NSFC (No.92270204), Youth Innovation Promotion Association CAS, Beijing Nova Program, and National Natural Science Foundation of China (No.62276149).
+Appendix
+Appendix A Dataset
+To ensure our dataset cover a sufficiently broad range of harmful topics, we conducted a comprehensive classification and statistical analysis.
+In terms of the question taxonomy in HarmBench [24], our dataset covers all 7 categories, with the respective distribution as follows: Cybercrime & Unauthorized Intrusion (16.7%), Chemical & Biological Weapons/Drugs (8.3%), Copyright Violations (10%), Misinformation & Disinformation (11.7%), Harassment & Bullying (10.8%), Illegal Activities (24.2%), and General Harm (18.3%).
+Due to the imbalance in the quantity of categories among the originally collected 100 questions from public datasets, with a significant lack of content related to Copyright Violations and Misinformation & Disinformation, our dataset expansion focused primarily on these two categories.
+Appendix B Open Source LLMs
+The detailed information about open source LLMs we used in our experiments is as follows:
+  • LLAMA-2-13B-Chat is LLAMA-2-13B fine-tuned with SFT and RLHF, it surpasses open-source chat models in helpfulness and safety, setting a robust baseline for further open-source LLM advancements.
+[…]
+  • Mistral-7B-Instruct is fine-tuned on public instruction datasets using SFT, it outperforms all preceding 7B models on instruction-following tasks, showcasing significant adaptability and performance.
+  • Mixtral8x7B-Instruct combines SFT with DPO to enhance instruction responsiveness and reduce biases, outperforming models like GPT-3.5 Turbo in human evaluations.
+Moreover, Mixtral features a Mixture of Experts architecture, setting it apart in terms of design and performance.
+  • Zephyr-7B utilizes distilled SFT and distilled DPO to align closely with user intent, setting new performance baselines for 7B models without the need for human annotation, and efficiently outperforming similar-sized models.
+Appendix C Dialogue Templates
 Upon reviewing the dialogue formatting procedures of open-source Large Language Models, it has been observed that they universally incorporate specific tokens to delineate the query from the completion.
-Examples of these formatting templates from some open-source models are provided, where the tokens used to separate the query and completion segments are highlighted for clarity.
-Dialog Template of Vicuna
-{{SYSTEM PROMPT}} USER: {{USER QUERY}} ASSISTANT: {{LLM COMPLETION}}
-Dialog Template of Mistral
-[INST] {{SYSTEM PROMPT}} {{USER QUERY}} [/INST] {{LLM COMPLETION}}
-Dialog Template of Zephyr
-<|system|> {{SYSTEM PROMPT}}</s> <|user|> {{USER QUERY}} </s> [<|assistant|>] {{LLM COMPLETION}}
-Dialog Template of Mixtral
-[INST] {{SYSTEM PROMPT}} {{USER QUERY}} [/INST] {{LLM COMPLETION}}
-Dialog Template of ChatGLM3
-<|system|> {{SYSTEM PROMPT}} <|user|> {{USER QUERY}} <|assistant|> {{LLM COMPLETION}}
-Dialog Template of Nous-Hermes-2-Mixtral-8x7B
-<|im_start|>system {{SYSTEM PROMPT}}<|im_end|> <|im_start|>user {{USER QUERY}}<|im_end|> <|im_start|>assistant {{LLM COMPLETION}}
-Appendix B Derivation of Log Perplexity Inequalities
+Examples of these dialog templates from open-source models are provided, where the separating tokens are highlighted for clarity.
+Dialog Template of Vicuna {{SYSTEM PROMPT}} USER: {{USER QUERY}} ASSISTANT: {{LLM COMPLETION}}
+Dialog Template of Mistral [INST] {{SYSTEM PROMPT}} {{USER QUERY}} [/INST] {{LLM COMPLETION}}
+Dialog Template of Zephyr <|system|> {{SYSTEM PROMPT}}</s> <|user|> {{USER QUERY}} </s> [<|assistant|>] {{LLM COMPLETION}}
+Dialog Template of Mixtral [INST] {{SYSTEM PROMPT}} {{USER QUERY}} [/INST] {{LLM COMPLETION}}
+Dialog Template of ChatGLM3 <|system|> {{SYSTEM PROMPT}} <|user|> {{USER QUERY}} <|assistant|> {{LLM COMPLETION}}
+Dialog Template of Nous-Hermes-2-Mixtral-8x7B <|im_start|>system {{SYSTEM PROMPT}}<|im_end|> <|im_start|>user {{USER QUERY}}<|im_end|> <|im_start|>assistant {{LLM COMPLETION}}
+Appendix D Derivation of Log Perplexity Inequalities
 Reflecting on the autoregressive nature of LLMs delineated in Section 2.1, the likelihood of an LLM generating a response $d$ given a context $x$ can be decomposed as follows:
-$\begin{split}\pi_{\Theta}(y=d|x)=\pi_{\Theta}(y_{1}=d_{1}|x)\prod_{i=1}^{m-1}{\pi_{\Theta}(y_{i+1}=d_{i+1}|x,...,y_{i}=d_{i})}\end{split}$
-By applying the logarithm to both sides of the equation and normalizing by the number of tokens in the response, we derive the equation:
-$\begin{split}&\frac{1}{m}\log\pi_{\Theta}(y=d|x)=\\ &\frac{1}{m}\left(\log\pi_{\Theta}(y_{1}=d_{1}|x)+\sum_{i=1}^{m-1}{\log\pi_{\Theta}(y_{i+1}=d_{i+1}|x,...,y_{i}=d_{i})}\right)\end{split}$
+$\begin{split}\small\pi_{\Theta}(y=d|x)=\pi_{\Theta}(y_{1}=d_{1}|x)\prod_{i=1}^{m-1}{\pi_{\Theta}(y_{i+1}=d_{i+1}|x,...,y_{i}=d_{i})}\end{split}$
+By applying the logarithm to both sides and normalizing by the number of tokens in the response, we derive the equation:
+$\small\begin{split}&\frac{1}{m}\log\pi_{\Theta}(y=d|x)=\\ &\frac{1}{m}\left(\log\pi_{\Theta}(y_{1}=d_{1}|x)+\sum_{i=1}^{m-1}{\log\pi_{\Theta}(y_{i+1}=d_{i+1}|x,...,y_{i}=d_{i})}\right)\end{split}$
 This expression is essentially the negation of the model’s log-perplexity for response $d$ given a context $x$.
 Correspondingly, according to inequalities 6 and the aforementioned equation, we have inequalities 10:
-$\begin{split}&\log PPL(y=d|x^{\prime})-\log PPL(y=d|x)>0,~{}\forall~{}d\in\mathcal{D}_{declination}\\ &\log PPL(y=d|x^{\prime})-\log PPL(y=d|x)<0,~{}\forall~{}d\in\mathcal{D}_{cooperation}\end{split}$
-Here, $PPL$ denotes the model’s perplexity, with $x$ representing the context where harmful content resides within the query, and $x^{\prime}$ signifying the context where harmful content presents in the completion.
-Appendix C System Prompts of Target Models
+$\small\left\{\begin{split}&\log PPL(y=d|x^{\prime})-\log PPL(y=d|x)>0,~{}\forall~{}d\in\mathcal{D}_{declination}\\ &\log PPL(y=d|x^{\prime})-\log PPL(y=d|x)<0,~{}\forall~{}d\in\mathcal{D}_{cooperation}\end{split}\right.$
+Here, $PPL$ denotes perplexity, $x$ is the context where harmful content resides within the query, and $x^{\prime}$ signifying the context where harmful content presents in the completion.
+Appendix E System Prompts of Target Models
 For the widely used large language models, we use their default system prompts as follows:
@@
 ChatGPT Series You are a helpful assistant.
+Appendix F LLAMA-2’s Update on System Prompt
+LLAMA2 updated its system prompt by removing the strict version, as it impaired the model’s normal functionality. The official updates are shown below:
+Observed Issue: We received feedback from the community on our prompt template and we are providing an update to reduce the false refusal rates seen.
+False refusals occur when the model incorrectly refuses to answer a question that it should, for example due to overly broad instructions to be cautious in how it provides responses.
+Updated approach: Based on evaluation and analysis, we recommend the removal of the system prompt as the default setting.
+Pull request #626 removes the system prompt as the default option, but still provides an example to help enable experimentation for those using it.